### v3

# The PLATO field selection process

11institutetext: INAF – Osservatorio Astronomico di Padova, vicolo dell’Osservatorio 5, 35122 Padova, Italy 22institutetext: Dipartimento di Fisica e Astronomia “Galileo Galilei”, Università degli Studi di Padova, Vicolo dell’Osservatorio 3, 35122 Padova, Italy 33institutetext: Deutsches Zentrum für Luft- und Raumfahrt (DLR), Institut für Planetenforschung, Rutherfordstraße 2, 12489 Berlin-Adlershof, Germany 44institutetext: INAF – Osservatorio Astrofisico di Catania, Via S. Sofia 78, 95123, Catania, Italy 55institutetext: INAF – Osservatorio Astronomico di Roma, Via Frascati, 33, 00078 Monte Porzio Catone (RM), Italy 66institutetext: SSDC-ASI, Via del Politecnico, snc, 00133 Roma, Italy 77institutetext: Institute of Astronomy, KU Leuven, Celestijnenlaan 200D, 3001, Leuven, Belgium 88institutetext: INAF–Osservatorio Astronomico di Palermo, Piazza del Parlamento, 1, I-90129, Palermo, Italy 99institutetext: Deutsches Zentrum für Luft- und Raumfahrt (DLR), Institut für Optische Sensorsysteme, Rutherfordstraße 2, 12489 Berlin-Adlershof, Germany 1010institutetext: Aix-Marseille Université, CNRS, CNES, Laboratoire d’Astrophysique de Marseille, Technopôle de Marseille-Etoile, 38, rue Frédéric Joliot-Curie, 13388 Marseille cedex 13, France $1$institutetext: Max-Planck-Institut für Sonnensystemforschung, Justus-von-Liebig-Weg 3, 37077 Göttingen, Germany 1212institutetext: LESIA, CNRS UMR 8109, Université Pierre et Marie Curie, Université Denis Diderot, Observatoire de Paris, 92195 Meudon, France1313institutetext: European Space Agency (ESA), European Space Research and Technology Centre (ESTEC), Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 1414institutetext: Centro de Astrobiología (CSIC–INTA), Depto. de Astrofísica, 28692 Villanueva de la Cañada, Madrid, Spain 1515institutetext: Center for Space and Habitability, University of Bern, Bern, Switzerland. 1616institutetext: Department of Physics, University of Warwick, Gibbet Hill Road, Coventry CV4 7AL, UK 1717institutetext: Armagh Observatory & Planetarium, College Hill, Armagh, BT61 9DG, UK 1818institutetext: Institute of Geological Sciences, Freie Universität Berlin, Malteserstraße 74-100, 12249 Berlin, Germany 1919institutetext: Institute of Astronomy, KU Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 2020institutetext: Observatoire de Genève, Université de Genève, Chemin Pegasi 51, 1290 Sauverny, Switzerland
The PLATO field selection process
II. Characterization of LOPS2, the first long-pointing field
[…]
(Submitted 20 September 2024 / Accepted 11 January 2025)
PLAnetary Transits and Oscillations of stars (PLATO) is an ESA M-class mission to be launched by the end of 2026 to discover and characterize transiting planets around bright and nearby stars, and in particular habitable rocky planets hosted by solar-like stars. Over the mission lifetime, an average of 8% of the science data rate will be allocated to Guest Observer programs selected by ESA through public calls. Hence, it is essential for the community to know in advance where the observing fields will be located. In a previous paper, we identified two preliminary long-pointing fields (LOPN1 and LOPS1) for PLATO, respectively in the northern and southern hemispheres. Here we present LOPS2, a slightly adjusted version of the southern field that has recently been selected by the PLATO Science Working Team as the first field to be observed by PLATO for at least two continuous years, following the scientific requirements. In this paper, we describe the astrophysical content of LOPS2 in detail, including known planetary systems, bright stars, variables, binary stars, star clusters, and synergies with other current and future facilities.
Key Words.:
Catalogues – Astronomical data bases – Techniques: photometric – Planetary systems – Planets and satellites: detection – Stars: clusters
[…]
1 Introduction
PLAnetary Transits and Oscillations of stars (PLATO; Rauer et al. 2024) is an ESA M-class mission primarily designed to detect a large number of planetary systems hosted by nearby and bright stars through the transit technique. Among many other aims, the goal of this mission is the discovery and confirmation of Earth-like planets, that is, habitable rocky planets hosted by solar-type dwarfs (Heller et al., 2022; Matuszewski et al., 2023). Very accurate stellar parameters, including stellar age at 10% accuracy, [^1] For solar-like stars. will be extracted from the light curves themselves through asteroseismological analysis (Cunha et al., 2021; Bétrisey et al., 2023). This will enable the discovered planetary systems to be placed in a consistent evolutionary context. Furthermore, follow-up and confirmation of candidate planets discovered by PLATO is part of the mission through the Ground-based Observation Program (GOP) of the PLATO Consortium. We refer the reader to Rauer et al. (2024) for a recent, detailed review of the mission.
During its nominal four-year mission, PLATO will continuously monitor one or two pre-selected fields for at least two years each during the so-called Long-duration Observation Phase (LOP). The mission has been designed to have the capability to observe additional fields for shorter intervals, at least two months each (Step-and-stare Observation Phase; SOP). The 24 “normal” cameras (NCAMs) of PLATO are not all co-aligned, but rather split into four groups with an angular offset (9.2) with respect to the satellite bore-sight (Pertenais et al., 2021). As a consequence, the overall field of view (FOV) of about 2 14921492\,1492 149 $\textrm{deg}^{2}$ is covered by a variable number of NCAMs, from six to 24, depending on the specific line of sight considered, as illustrated by Fig. 1. In terms of sky area, approximately 325 deg2 will be covered by 24 NCAMs, 153 deg2 by 18 NCAMs, 847 deg2 by 12 NCAMs, and 824 deg2 by six NCAMs (assuming a corrected FOV of $19.2^{\circ}$ and CCD gaps of 1.3 mm; Pertenais et al. 2021), although the exact numbers could slightly change according to the actual optical performances once in flight.
Due to telemetry constraints, PLATO will not be able to download full-frame images at high cadence. [^2] Full-frame images will be acquired during the commissioning phase, at least once every 90 days. Instead, it will perform a significant part of the photometric analysis onboard. Only light curves, centroid data, and a number of stamp-like “imagettes” will be transmitted back to Earth, meaning that targets have to be selected in advance. PLATO targets are divided into four main samples, summarized in Table 1, with different requirements on magnitude, spectral type, and noise-to-signal ratio (NSR); three of such samples are composed by main-sequence or subgiant stars with a spectral type from F5 to K7 (P1, P2, P5), while a fourth one contains only M dwarfs (P4). The PLATO Scientific Requirement Document [^3] ESA reference: PTO-EST-SCI-RS-0150. (SciRD) identifies P1 ($V<11$, $\textrm{NSR}<50$ ppm in one hour) as the highest-priority sample, and sets a requirement of at least 15 0001500015\,00015 000 P1 targets to be observed during the LOPs. We refer to the SciRD, Nascimbeni et al. (2022), and Montalto et al. (2021, hereafter quoted as N22 and M21) for a complete summary of the definitions and requirements on the PLATO samples.
Figure 1: Geometry of the PLATO field FOV. In this figure, the field is centered at the origin (0,0) of a generic spherical reference frame (units are degrees; projection is orthographic). The number of “normal” cameras covering a given line of sight is color coded. The four blue shades, from dark to light, map the regions covering 325/153/847/824 deg2 observed respectively with 24, 18, 12, six cameras (and corresponding to four, three, two, and one group of six co-pointing telescopes each).
PLATO is currently on schedule to be launched at the end of 2026 and to start its routine scientific operations in mid-2027. Most raw and calibrated data products from PLATO will be made publicly available as soon as possible. [^4] With the exception of GO programs and reserved targets allocated to the Plato Mission Consortium, within their proprietary time. The astronomical community has the opportunity to be deeply involved at different stages: preparation, data analysis, and follow-up. PLATO data products will be released quarterly. Each set of data from a three-month observation period will be made available approximately six months later. For around 20 0002000020\,00020 000 bright targets prioritized for ground-based follow-up observations, the data products will be delivered roughly 1.2 years after the corresponding three-month observation period. On top of that, over the mission lifetime, an average of 8% of the PLATO science data rate will be allocated to Guest Observer (GO) programs selected by ESA through open calls, so it is essential to select and communicate each field that is going to be pointed at with a proper time margin.
In a previous work (Paper I; N22), we described all the steps of the process that led to the identification of two provisional fields (labeled LOPN1 and LOPS1) as candidates for the LOP. Back then, it was still unknown which field would be selected as the first one. Also, it was anticipated that the final LOP fields could have had small changes in position and/or rotation angle. In this paper, we present LOPS2, the long-pointing field finally chosen by the ESA PLATO Science Working Team (PSWT) as the first field to be observed by PLATO. After summarizing the fine-tuning process and the general properties of LOPS2 in Section 2, we review and discuss its astrophysical content in Section 3. Finally, the conclusions and some prospects for the future work to be done are given in Section 4. A glossary of the most common acronyms used throughout this work is compiled in Table 3 in the Appendix.
2 The choice of the first long-pointing field
Paper I gave a very detailed description of the process that led to the selection of LOPS1 and LOPN1 as candidate long-duration fields for PLATO. In order to fine-tune their position/rotation angle and to help the SWT to select the first LOP, three main criteria were considered:
  • Duty cycle: The PLATO attitude is defined by the pointing direction and the rotation around the mean line of sight. Because of the geometry of the overlapping cameras, different rotation angles result in different sets of stars observed with a given photometric precision (with a different number of cameras). For pointing directions with ecliptic latitudes between $63^{\circ}<|\beta|<70^{\circ}$, the choice of the rotation angle around the mean line of sight is constrained by the geometry of the spacecraft. There is always a possible rotation angle that meets the duty cycle requirements (93% of time on-target), but some angles result in incompatible attitudes. For pointing directions with $|\beta|>70^{\circ}$ all choices of the rotation angle are compatible with the duty cycle requirements.
  • Rotation angle: the choice of a different rotation angle should be investigated for any technical or scientific advantage.
  • Synergies: the field with the most interesting synergies with other facilities should be given higher priority, in particular with regard to the radial-velocity follow-up process, considering the time frame of the mission.
In what follows, we discuss the impact of these three key ingredients on the final decision. The final decision is presented and discussed in Section 2.4.
2.1 Duty cycle
As anticipated in Paper I, Section 2.4, the need of keeping the solar panels of PLATO at a nominal level of solar irradiation while constantly staring at the same LOP field for $\geq 2$ years implies that the roll angle of the spacecraft must be moved at regular intervals throughout the year. Since the PLATO field, by design, shows a $90^{\circ}$ symmetry, it is natural to perform $90^{\circ}$ rotations each quarter, also known as quarterly rolls. A similar observing strategy was already devised for the original Kepler mission (Borucki et al., 2010). The ecliptic latitude $\beta$ is the crucial parameter to be investigated, as it determines the incidence angle of the solar rays on the panels. There is a critical value of $|\beta|$ above which an observing quarter can start at any moment, without any need of interruptions before the first roll.
[…]
  • the prioritization metric as defined in Paper I is 1% smaller (0.980 vs. 0.990), that is, the difference is negligible;
  • the number of P1 targets in LOPS2, defined as in Table 1 and evaluated with the most updated version of the PLATO Input Catalog (PIC) (v.2.0.0), is 0.4% lower, but still well within the margins of SciRD requirements.
In other words, LOPS2 appears as an optimal alternative, having properties very similar to LOPS1, but allowing us to be much more efficient during mission operations. For example, the in-orbit commissioning will take place before the start of the nominal science operations. In order to avoid constraints on the choice of the rotation angle for commissioning (and hence in the duty cycle), the pointing direction for the commissioning field must have $|\beta|>70^{\circ}$, which was an incompatible constraint with LOPS1. We therefore consider LOPN1 (as defined in Paper I) and LOPS2 (as defined above; detailed coordinates in Table 2) as candidate LOP fields hereafter.
Table 1: Definitions of the main PLATO stellar samples.
sample SpT lum. class mag. limit noise
P1 F5-K7 dwarfs and subgiants $V<11$ $<50$ ppm
P2 F5-K7 dwarfs and subgiants $V<8.5$ $<50$ ppm
P4 M dwarfs $V<16$
P5 F5-K7 dwarfs and subgiants $V<13$
 [^5] The rows give: the name of the PLATO stellar sample, the spectral type, the luminosity class, the limiting magnitude in the $V$ band, and the noise limit in ppm in one hour (see also Section 3.1).
Although it has a negligible impact on the duty cycle, it is worth reminding that the the dates of the quarterly rotations along the orbit are uniquely determined by the choice of the attitude. This is a consequence of the duty cycle requirement and the geometry of the spacecraft. For LOPS2 and LOPN1, the quarterly rotations will take place at the end of January, April, July, and October (the exact dates differ by few days in both hemispheres). For a scheduled launch end 2026 and considering that the commissioning shall take less than 90 days, it could very well be that the first quarter is significantly shorter than the nominal duration of approximately 90 days, similar to the situation experienced by CoRoT and Kepler.
2.2 Rotation angle
In paper I, both fields were presented at zero rotation angle, which, in our convention, implies that one side of the field is parallel to the galactic plane. We investigated the impact of the rotation angle $\varphi$ by generating nine simulated catalogs for both LOPN1 and LOPS2 by varying the angle by $10^{\circ}$ intervals (from $\varphi=0^{\circ}$ to $80^{\circ}$). For the resulting 18 catalogs, extracted from the PIC 2.0.0, we added a NSR column (calculated with PINE; Börner et al. 2024), in order to estimate the number of P1-P2-P4-P5 counts and to assess the visibility of our main targets of interest.
As expected, given the large size of the PLATO field and its fair degree of radial symmetry, $\varphi$ has a mostly negligible impact on the number of targets. For both LOPN1 and LOPS2 the P1 counts are constant within $\sim 1\%$, with the highest number corresponding to the $\varphi=0^{\circ}$ case (Fig. 14, left panel). On the other hand, P4 and P5 counts are virtually unaffected, showing just $<0.5\%$ variations, that is at the level of the predicted Poissonian noise (Fig. 14, right panel). We can conclude that $\varphi$ has no practical effects on the planet yield of PLATO.
It is worth considering whether there is any other reason to rotate the field, for instance due to specific astrophysical objects we would like to include, or avoid, in the LOP field. As for the former, we verified that 1) the number of known planetary systems is not significantly affected by $\varphi$ for LOPN1, and is actually maximized at $\varphi=0^{\circ}$-$10^{\circ}$ for LOPS2 due to an overdensity of TESS-discovered planets lying in the galactic southwest corner of the field (as we show in Section 3.4); and 2) no open cluster among the ones of potential interest for PLATO (i. e., nearby, sparse, and mature) slips in or out of the field for any rotation angle choice, as it is shown in Section 3 for LOPS2. More in general, the range of galactic latitudes probed is always $0^{\circ}<|b|\lesssim 50^{\circ}$ regardless of $\varphi$, and therefore any property correlated with $b$, such as the amount of stellar crowding and the diversity of stellar populations probed (in terms of age, kinematics, metallicity, etc.) is approximately the same.
Finally, it makes sense to investigate whether there is any way to rotate the field in order to move any extremely bright star ($V<1$) off the silicon. Paper I identified Deneb and Vega as the brightest sources on LOPN1, and Sirius and Canopus for LOPS1. Moving from LOPS1 to LOPS2 also implies that Sirius is now left outside the focal plane for any choice of $\varphi$, while there is no consequence on the other three stars, that are unavoidable and always monitored with the same number of NCAMs (18-24 for Vega, 6 for Deneb, 24 for Canopus) for almost any possible rotation angle. [^6] It could be possible in principle to fit Deneb within a $\sim 30^{\prime}$ CCD gap of the northern field at $\varphi\simeq 11^{\circ}$, but the technical feasibility of such a choice and its possible unintended consequences (see for instance Porterfield et al. 2016) need to be assessed. This will be done at a further stage, should LOPN1 be scheduled for observation. Fig. 15 (but also Section 3.2 below) investigates this point a bit further, looking at how the number of bright stars (up to $V<3$) in LOPS2 changes as a function of $\varphi$, at $10^{\circ}$ intervals from $0^{\circ}$ to $80^{\circ}$ (the symmetry of the PLATO field makes it invariant after $90^{\circ}$ rotations). The nominal $\varphi=0^{\circ}$ case is the only one for which all the five $21<V<$ 2 stars ($\delta$ CMa, $\epsilon$ CMa, $\gamma_{2}$ Vel, $\delta$ Vel, $\epsilon$ Car) always land on 6 NCAMs regions, while for instance four of them fall on 12 NCAMs in the $\varphi=40$-$50^{\circ}$ case. Finally, also the distribution of $32<V<$ 3 stars does not change significantly as a function of $\varphi$.
In summary, we conclude that there is no reason to change the rotation angle of both LOPN1 and LOPS2 from its initial value of $\varphi=0^{\circ}$. We implicitly assume zero rotation (i.e., one side of the field almost tangent to the galactic plane) hereafter, as plotted in the all-sky projection in Fig. 2.
Figure 2: All-sky Aitoff projection in galactic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively. This sky chart is also plotted in equatorial and Ecliptic coordinates in Fig. 17-18.
2.3 Synergies and follow-up
The last possible criterion remaining to guide the choice between LOPN1 and LOPS2 are the synergies with other past, present and forthcoming observing facilities, including (and with a particular attention to) those to be involved with the PLATO follow-up process during and after the nominal scientific observations of the first field, approx. from mid-2027 with a minimum duration of two years.
As for the previous space-based transit surveys, we first remind that no overlap between a LOP field and K2 or CoRoT (Howell et al., 2014; Auvergne et al., 2009) is possible at all, since the latter fields are all located close to the ecliptic by design, inaccessible during the LOP phase (Fig. 2). The TESS continuous viewing zone (CVZ) at $|\beta|>78^{\circ}$, on the other hand, is monitored at 100% in LOPN1 and $\gtrsim 90\%$ in LOPS2, mostly in six- and 12-camera regions. Finally, the Kepler field is fully included in the LOPN1, almost entirely with at least 12 NCAMs and up to 24. While the scientific case of following-up the Kepler TTV systems with PLATO is very solid (Jontof-Hutter et al., 2021a, b), we emphasize that it should not be seen as a main driver of the mission, but rather as a potential source of interesting additional science. [^7] It is worth mentioning that the follow-up of the Kepler field is part of the science goals of the proposed Earth-2.0 mission (Ge et al., 2022). Indeed, the last word is on the follow-up availability and timeliness, since the confirmation of the PLATO candidates and eventual mass determination is part of the mission products (and goals) through the Ground-based Observation Program (GOP) of the PLATO Consortium.
We already noted in Paper I that the range of declination spanned by any compliant LOP field (including LOPN1 and LOPS2) hinders the follow-up of a large fraction of the northern field with most southern facilities and vice versa, with the exception of a strip in the low-$|\delta|$ region of the fields, whose size depends on the maximum acceptable airmass $X$ and the geographical latitude $\psi$ of the observing site. Such a strip can be typically monitored from the opposite hemisphere only for 1-2 months per year at reasonably low airmass values ($2.01.5\lesssim X<$ 2.0; see Paper I), making ultra-high precision ($<1$ m/s) RV measurements unfeasible or exceedingly difficult for the vast majority of P1 targets. A detailed analysis on how the visibility of LOPS2 and its sub-regions covered by six, 12, 18, 24 NCAMs is dependent on $\psi$ and $X$ is shown in Fig. 16. As a reference value, we note that at the Observatorio Roque de Los Muchachos (ORM) only 16% of LOPS2 is reachable at $X<2$ (and only in the 6-12 NCAMs region), while at Mauna Kea 36% of LOPS2 can be observed with the same requirement (but only 8% of the 24 NCAMs region can be reached).
A survey of the high-precision RV spectrographs available to the European community reveals that most of them are located in Chile: they include the already operational HARPS (Mayor et al., 2003) and ESPRESSO (Pepe et al., 2021); and the planned ANDES, whose first light at the ESO Extremely Large Telescope (ELT) is currently foreseen in 2031 (Marconi et al., 2022). More in general, large-scale facilities such as ELT will also host other instruments with a huge potential in the follow-up process, such as coronagraphs and infrared spectrographs. Conversely, no large telescope of the 30-m class is currently at the construction stage in any northern observatory.
It is worth mentioning two other important survey projects in the south that could have interesting implications for the PLATO preparation and/or follow-up: LSST (Ivezić et al., 2019) at the Vera C. Rubin Observatory, which will map the whole southern hemisphere (including LOPS2) beginning in 2025, and 4MOST at the VISTA telescope (de Jong et al., 2012), starting its consortium surveys also in 2024. The southern hemisphere will also host the Giant Magellan Telescope (GMT; Johns et al. 2012), which is scheduled to be operational in the early 2030s.
2.4 LOPS2 as first LOP field
On the basis of what has been discussed in the previous Section, in June 2023 the SWT has formally approved LOPS2 (at $\varphi=0^{\circ}$) as the first LOP field. A full-sky map in galactic coordinates including LOPS2 and LOPN1 is presented in Fig. 2; the same full-sky map is also plotted in equatorial and Ecliptic coordinates in Fig. 17-18; a zoomed-in chart centered on LOPS2 is plotted in Fig. 3 in both galactic and equatorial coordinates. All the relevant quantities for LOPS2 are listed in Table 2, including the number of targets for each PLATO sample. The latter are calculated by adopting PIC 2.0.0 as input catalog and conservatively assuming an end-of-life scenario with 22 surviving NCAMs (“EOL 22”). We refer the reader to Section A in the Appendix on how to check if a given object falls within the LOPS2 boundaries.
Table 2: Some properties of the LOPS2 field.
parameter value notes
$\alpha$ [deg] 95.31043 ICRS
$\alpha$ [hms] 06:21:14.5 ICRS
$\delta$ [deg] --47.88693 ICRS
$\delta$ [dms] --47:53:13 ICRS
$l$ [deg] 255.9375 IAU 1958
$b$ [deg] --24.62432 IAU 1958
$\lambda$ [deg] 101.05940 Ecliptic
$\beta$ [deg] --71.12242 Ecliptic
P1 targets 8 235 SciRD req. 7 500
P2 targets 699 SciRD req. 500
P4 targets 12 415 SciRD req. 2 500
P5 targets 167 149 SciRD req. 122 500
 [^8] The rows give the coordinates for the field center in equatorial, galactic, and ecliptic reference frames and the number of targets available in the P1-P2-P4-P5 samples. The latter are calculated from PIC 2.0.0 assuming the “EOL 22” scenario (see Section 2.4).
Figure 3: LOPS2 footprint. Left panel: Orthographic projection in galactic coordinates centered on LOPS2. The field is color coded according to the number of NCAMs as in Fig. 1. The yellow circle at $\beta=-78^{\circ}$ approximates the southern TESS continuous viewing zone (CVZ), while the magenta circle is centered on the Large Magellanic cloud. Right panel: Same as in left panel but in equatorial coordinates.
[…]
The average interstellar extinction coefficient $A_{V}$ and its spatial dependence is shown in Fig. 4 for both the P1+P2 and P5 samples of LOPS2. In the former case we always got $A_{V}<0.07$; in the latter sample, a clear structure with $A_{V}\simeq 0.3$ appears, corresponding to some parts of the Vela-Puppis region (Cantat-Gaudin et al., 2019). Clearly, the main reason for the larger extinction of the P5 sample is the much larger average distance of its stars, due to a much fainter magnitude limit ($V<13$ for the P5 sample, $V<11$ for P1). No significant excess of extinction can be seen on these maps in the LMC region, because even our faintest and intrinsically more luminous targets lie always within 3 kpc from the Sun: the LMC is just a background object for the PLATO planet-hunting survey. See also Section 3.1 for some details on the distance distribution.
Figure 4: Interstellar extinction properties of LOPS2. Left panel: Average $V$-band extinction $A_{V}$ of the P1+P2 targets from PIC 2.0.0, calculated over HEALPix level-5 cells and represented in a color scale. Right panel: Same but for the P5 sample and over a finer level-6 grid. The high-extinction bump at $A_{V}\simeq 0.3$ is due to the closest parts of the Vela-Puppis region (Cantat-Gaudin et al., 2019).
3 Astrophysical content of LOPS2
3.1 Main PLATO targets
Adopting PIC v. 2.0.0 as input catalog, and calculating the NSR through the PINE code (Börner et al., 2024), we can calculate the counts for the four main samples of PLATO (defined according to the formal requirements of the SciRD):
  • 8 23582358\,2358 235 P1 targets, i. e., FGK dwarfs and subgiants with $\textrm{NSR}\leq 50$ ppm in one hour and $V\leq 11$;
  • $699$ P2 targets, i. e., FGK dwarfs and subgiants with $\textrm{NSR}\leq 50$ ppm in one hour and $V\leq 8.5$;
  • 12 4151241512\,41512 415 P4 targets, i. e., M dwarfs with $V\leq 16$;
  • 167 149167149167\,149167 149 P5 targets, i. e., FGK dwarfs and subgiants with $V\leq 13$.
The definitions listed above are also summarized in Table 1. When the samples are computed as disjoint sets (that is, P2 is excluded from P1, and P1+P2 is excluded from P5), in LOPS2 we have 699 P2 targets, 7 53675367\,5367 536 P1 targets and 158 914158914158\,914158 914 P5 targets. The total sample of available targets in LOPS2 is therefore 179 564179564179\,564179 564 FGKM stars. We note that, with some exceptions and unless otherwise stated, targets in LOPS2 that are not included in the PLATO samples will not be observed by default and should be proposed in the GO program. The detailed policy will be communicated in advance of the first GO call. According to the ESA Science Management Plan (available on the ESA website, but also summarized in Rauer et al. 2024, Section 10) the first public release of the PIC, which also includes flags for P1-P2-P4-P5, will be released to the community nine months before the launch, concurrently with the call for the GO proposals. According to the current mission schedule, this is expected to happen in spring 2026.
The distances of individual solar-type LOPS2 targets (i.e., samples P1+P2 and P5) are plotted in Fig. 5, while the statistical distribution in terms of distance, interstellar extinction $A_{V}$ and magnitude is shown in the histograms of Fig. 19-20, for both the whole PIC and for a few subsamples of interest. Unsurprisingly for a magnitude-limited sample, and as also shown in Paper I, late-type, main-sequence targets such as G/K dwarfs are on average much closer to us with respect to early-type stars and (sub-)giants, due to their fainter absolute magnitude. As a consequence, the former have also lower interstellar extinction values with respect to the latter (right panel of Fig. 19).
The histograms of some other astrophysical properties of the P1+P2 and P5 samples are plotted in Fig. 21-22. The spatial distributions of the P1, P2, P4, P5 samples are plotted in the four panels of Fig. 23: P2 and P4 stars are almost perfectly homogeneously distributed, because they both lie relatively close to the Sun, and are purely magnitude-limited samples. [^9] Formally, the P2 definition also includes a $\textrm{NSR}\leq 50$ ppm requirement but with the current design it is always satisfied even at its $V=8.5$ faint limit. P5 is also magnitude-limited, but being on average much fainter is also farther from the Sun, and the consequent line-of-sight effects due to the galactic disk are revealed by the significant density gradient along galactic latitude. Finally, P1 is concentrated toward the inner region of LOPS2, observed with 18/24 NCAMs, where the noise requirement is easier to be met at equal magnitude.
Figure 5: Distance of the P1/P2 and P5 samples in LOPS2. Left panel: Distribution of the 8 23582358\,2358 235 P1+P2 stars in LOPS2 in the ($R_{\star}$, distance) plane. The stellar effective temperature from the PIC v2.0.0 is color coded. Right panel: Same but for the P5 sample (167 149167149167\,149167 149 stars; P1/P2 not included). The distribution moves to larger distances due to the fainter magnitude limit of P5 ($V<13$ vs. $V<11$).
3.2 Bright stars
Figure 6: Bright stars and stellar crowding in LOPS2. Left panel: Orthographic projection in galactic coordinates centered on LOPS2. The field is color coded according to the number of NCAMs as in Fig. 1. The brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3), gray ($43<V<$ 4). Right panel: Same but the footprint of LOPS2 is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The color scale increases proportionally to density from orange to blue with a transition at $\sim 400$ targets per $\textrm{deg}^{2}$, at $b\simeq-15^{\circ}$.
As part of the field selection process, and as anticipated in Section 2.2, we investigated the presence of very bright stars on LOPS2. We used the Yale Bright Star Catalogue version 5 revised (YBSC; Hoffleit & Warren 1995) as input, being a reasonably complete and accurate census of all the stars brighter than $V\simeq 6.5$. The strategy of PLATO for moderately saturated stars imaged on the NCAMs (i.e., between magnitudes 4 and 8 in the PLATO Vegamag system, and approximately [^10] The exact saturation limit is largely determined by factors other than the color term $P-V$, including the position on the FOV, the PSF shape, the actual in-orbit CCD gain and optical transmission, etc. also in the $4\lesssim V\lesssim 8$ range) is to use larger imagettes (saturated star windows) that extend up to $256\times 6$ pixels in size to extract the photometry, if the additional telemetry cost needed to download more data is justified. For stars brighter than magnitude 4, a different strategy is being discussed. For example, smearing photometry is technically possible (White et al., 2017; Pope et al., 2019), but alternatives are being considered.
In any case, knowing the position of such bright sources could be of interest since their bright halos, bleeding columns and smearing trails can potentially contaminate any other fainter target in their neighborhood. A discussion on the technical issues related to CCD saturation, blooming, non-linearity as well as optical ghosting is discussed in Verhoeve et al. (2016); Pertenais et al. (2022); Jannsen et al. (2024a) and will be better explored in the forthcoming papers of this series.
Among the 712 YBSC stars lying in the LOPS2, only 47 are brighter than $V=4$ (Fig. 6, left plot). With the only exception of Canopus = $\alpha$ Car (also by far the brightest one; $V=-0.72$), Phact = $\alpha$ Col, and $\tau$ Pup, all the 12 extremely bright stars at $V<3$ are located close to the galactic plane at $|b|<15^{\circ}$, where also the average stellar density increases very rapidly as the line of sight crosses the galactic disk (Fig. 6, right plot). Not surprisingly, all of them are early-type and/or evolved giants, so none would fit the spectral type requirement to be included in the PLATO P2 sample.
Scrolling down the list, $\zeta$ Dor A = HD 33262 at $V=4.68$ is the brightest star in LOPS2 to be included in the PIC: it is the F7V component of a very wide visual binary. It is also known as a young and active star surrounded by a debris disk (Dodson-Robinson et al., 2011). B Car = HD 68456 = HR 3220 comes as a close second at $V\simeq 4.73$; it is a single-lined spectroscopic binary whose F7V primary was also identified as a field blue straggler by Fuhrmann et al. (2011), transferring mass to its low-mass white dwarf companion on a $P=899$ d orbit. 212 Pup = HD 64379, at $V=5.05$, is the brightest binary system made of main-sequence stars (F7V+K5V) in our sample, with a narrow separation ($3"$). Finally, the brightest single solar twin in the P2 sample is HD 59967 ($V=6.66$), a young G2V star (Lorenzo-Oliveira et al., 2018) with a debris disk (Pearce et al., 2022), showing evident high levels of stellar activity and with no detected planetary companions from FEROS (Zakhozhay et al., 2022) and HARPS (Grandjean et al., 2023).
3.3 Nearby stars
Among the 562 known stars within 10 pc from the Sun (Reylé et al. 2021, 2023-02-06 version), 14 are located within LOPS2. Only seven of them are bright enough ($V\lesssim 16$) to be reliably monitored by PLATO. [^11] The excluded targets are all at $G\gtrsim 20$, that is, way too faint for PLATO to deliver reliable photometry. Six of them are M dwarfs or sub-dwarfs, spanning spectral types from sdM1V to M4.5V; sorted by increasing $G$ magnitude and spectral type: GJ 191, CD-44 3045 B and A, AP Columbae, L 230-188, SCR J0740-4257. The seventh one is LAWD 26 = GJ 293, a bright hydrogen-rich white dwarf (WD; DA; $G\sim 13.7$, $V\sim 14.1$), spectroscopically confirmed by Bell & Rodgers (1964) and recently identified by Sanderson et al. (2022) as one of the most promising white dwarf targets for the discovery of astrometric planets with Gaia.
Other nearby and reasonably bright WDs can be identified from the catalog by Gentile Fusillo et al. (2021): 86 targets with $G<15$ lie within the LOPS2 footprint. [^12] We emphasize that WDs will not be included in the main PIC subsamples (P1-P2-P4-P5) due to the spectral class requirement (Section 3.1), unless they are flagged as planetary hosts or candidate hosts (See Section 3.4). Within the technical limits, though, they can be included as “scientific calibration and validation” targets (scvPIC; Zwintz 2024) or as GO targets. Among these, and besides the already mentioned LAWD 26, four additional stars show a parallax $\pi>50$ mas, that is, they are closer than 20 pc: two of them were also spectroscopically confirmed by O’Brien et al. (2023), both as DA spectral types.
It is worth mentioning that Kapteyn’s star = GJ 191 ($V\sim 8.8$, $K\sim 5.0$) is an old M1 sub-dwarf at only 3.9 pc from the Sun, and also the nearest known halo star (Kotoneva et al., 2005). It will be observed by PLATO through 12 cameras, and with an average NSR estimated around 18.4 ppm in 1 h according to the current noise budget calculations. Two planetary candidates were claimed to be discovered through RVs around GJ 191 by Anglada-Escude et al. (2014), but were later disproved as due to stellar activity by Robertson et al. (2015) and Bortle et al. (2021), who also noted that the star is photometrically very stable. As is discussed in Section 3.4, a coordinated RV follow-up of this target could take advantage of the simultaneous PLATO coverage to disentangle the contributions of stellar activity and the Keplerian signal from planets, if any.
CD-44 3045 A and B, also known as GJ 257, represent an equal-mass visual binary (M3V + M3V) with an angular separation of $2.4"$ at 2016.0 according to Gaia DR3. Hence, they will fall on the same $\sim 15"$ pixel of PLATO, and observed as a single object in combined light within the P4 sample.
3.4 Known planetary systems
It is of primary interest to identify all the already known planets and planetary candidates within LOPS2, since 1) it makes the vetting and follow-up process much more efficient, 2) there are many science cases for which a long-term and/or extreme-precision follow-up of already known planets can yield compelling scientific results (see for instance Jontof-Hutter et al. 2021b), and 3) the unprecedented photometric accuracy of PLATO may allow us to discover additional transiting planets around planet-hosting stars and/or confirm low-SNR candidates and mono-transits (Magliano et al., 2024).
For the reasons mentioned above, it has been decided that the final PIC will include all confirmed and candidate exoplanets known at the compilation date, so all the planetary systems mentioned in the following sections, if technically feasible with PLATO, will be forced into the catalog regardless whether they meet the P1-P2-P4-P5 SciRD requirement or not. We mention that a preliminary review of the known planetary systems located in LOPS2 was also presented by Eschen et al. (2024), although limited to transiting planets only and with a particular focus on the PLATO vs. TESS synergy. We note that their overall numbers are roughly consistent with ours, but a direct comparison is not possible because a combination of different catalogs and/or unreported catalog versions were employed in their match.
3.4.1 Transiting planets
We retrieved the list of known planets from Exo-MerCat (Alei et al. 2020, Alei et al. in prep.), a meta-catalog built by comparing and merging with an automated process scanning three input catalogs: Exoplanet Encyclopedia, [^13] https://exoplanet.eu/home/ the NASA Exoplanet Archive [^14] https://exoplanetarchive.ipac.caltech.edu/index.html and the Open Exoplanet Catalog. [^15] https://www.openexoplanetcatalogue.com/ In this Section we focus on the transiting sub-sample only, which is the most important one given the scientific opportunity they represent for PLATO.
Overall, 108 confirmed transiting planets are located in LOPS2 (Fig. 7, left panel and Table LABEL:table:transiting_planets); they belong to 84 distinct planetary systems. Among the 13 systems with multiple transiting planets there are five doublets (GJ 143, TOI-216, TOI-286, TOI-431, TOI-2525), six triplets (TOI-270, TOI-451, TOI-712, LHS 1678, L98-59161616L98-59 and GJ 143 lie very close to a LOPS2 external border (Fig. 7, left plot), therefore their inclusion in the target list (and the delivered photometric precision) will depend on the actual pointing accuracy and system performance. and HD 28109), and two high-multiplicity systems that look as perfect showcases for the photometric performances of PLATO:
  • TOI-700, a four-planet system hosted by a relatively bright ($V\simeq 13$) M2V star, including two super-Earths in or close to the habitable zone (Gilbert et al., 2020, 2023);
  • HD 23472 = TOI-174, a packed five-planet system (Trifonov et al., 2019; Teske et al., 2021) orbiting around a $V\simeq 9.7$ K dwarf, made of three confirmed super-Earths (with at least a suspected 5:3 mean-motion resonance) and two candidate high-density “super-Mercuries”, among the least massive planets to ever have been detected through RVs (Barros et al., 2022).
Among the mentioned multiple transiting systems we also identify three systems hosted by relatively young stars ($<10^{9}$ years): TOI-451, belonging to the Pisces Eridanus stream (120 Myr; Newton et al. 2021), and the field stars TOI-201 (870 Myr; Hobson et al. 2021) and TOI-712 (830 Myr; Vach et al. 2022).
There are also five additional entries in Exo-MerCat whose estimated masses fall within the brown dwarf (BD) regime (HIP 33609b, TOI-569b, TOI-811b, TOI-1496b) or across the planet/BD transition (HATS-70b; Zhou et al. 2019). They are plotted as yellow diamonds in Fig. 7.
Figure 7: Known planetary systems with at least one transiting planet in LOPS2 (orthographic projection in galactic coordinates). Left panel: Entries from the Exo-MerCat database (Section 3.4): single transiting planets (red circles), systems with multiple transiting planets (orange circles with labels; the number inside the circles is the multiplicity), transiting brown dwarfs (yellow diamonds). Right panel: Same projection but TESS candidates from the TOI database are plotted with yellow circles. Systems with multiple transiting candidates are marked with red squares (doublets) or magenta squares (triplets).
Figure 8: Known planetary systems with at least one transiting planet in LOPS2. Left panel: Period-radius diagram with the 108 transiting planets identified from Exo-MerCat and plotted in the left panel of Fig. 7. The “Neptunian desert” region as identified by Mazeh et al. (2016) and with the modified lower boundary by Deeg et al. (2023) is plotted in light gray. Planets belonging to multiple systems are marked with a gray square. Right panel: Period-radius diagram with the 612 candidate transiting planets from the TESS TOI database and yet unpublished, plotted in the right panel of Fig. 7
A closer look to the planetary parameters of the 108 transiting planets (Fig. 8, left panel) reveals a vast diversity of planetary properties. Adopting some common definitions from the literature, we can split our sample into four main classes:
  • 47 hot Jupiters (HJ; $R_{\mathrm{p}}>6$ $R_{\oplus}$, $P<10$ d). This sub-sample is unlikely to increase significantly in the future, given the almost perfect sensitivity of TESS and the existing ground-based surveys in this region of the parameter space. Notably, there are three ultra-hot Jupiters (UHJs; $T_{\mathrm{eq}}>2200$ K): WASP-121b (Delrez et al., 2016), KELT-25b (Rodríguez Martínez et al., 2020), and the already mentioned HATS-70b. WASP-121b is of particular interest since it spawned an impressively vast literature on its atmospheric characterization, including the study of its optical phase curve (Bourrier et al., 2020; Daylan et al., 2021) and claims of temporal variability as well (Wilson et al., 2021);
  • 14 warm or cool Jupiters (WJ/CJ); $R_{\mathrm{p}}>6$ $R_{\oplus}$, $P>10$ d), eight of them on orbital periods longer than 50 d. The longest period planet of our whole LOPS2 sample belongs to this sample: TOI-4562b, a temperate Jupiter analog on a $P\simeq 225$ d orbit (Heitzmann et al., 2023) recently claimed to show TTVs by Fermiano et al. (2024);
  • 24 Neptunians and mini-Neptunes ($1.7$ $R_{\oplus}<R_{\mathrm{p}}<6$ $R_{\oplus}$), including at least one planet undoubtedly within the “hot Neptune” desert (Mazeh et al. 2016; gray area in Fig. 8): NGTS-4b (West et al., 2019) plus two other more marginally so (TOI-451b, TOI-269b);
  • 23 rocky planets ($R_{\mathrm{p}}<1.7R_{\oplus}$), including four ultra short-period ones (USP; $P\lesssim 1$ d): TOI-206b ($P=0.74$ d; Giacalone et al. 2022), LHS 1678b ($P=0.86$ d; Silverstein et al. 2024), TOI-431b ($P=0.49$ d; Osborn et al. 2021) and TOI-500b ($P=0.55$ d; Serrano et al. 2022). The latter will be observed close to the center of LOPS2 through 24 NCAMs, so PLATO will be able to gather at least $\approx 1300$ full phase curves at $\textrm{NSR}\simeq 33$ ppm in one hour.
We calculated for each entry of the Exo-MerCat sample the so called ephemeris drift $\sigma(T_{0})$ at epoch 2027.0, that is, the expected 1-$\sigma$ error on the transit prediction (according to the most accurate ephemeris available) at the approximated epoch when PLATO is supposed to start its scientific operations. The median is $\sigma(T_{0})\sim 7$ min. Only ten planets show a drift larger than two hours; not surprisingly, they are mostly small and/or long-period planets observed by TESS on a limited number of sectors and not accessible by ground-based facilities. At $\sigma(T_{0})\simeq 1.8$ d, an extreme case is TOI-1338b = BEBOP-1b (the first circumbinary planet discovered by TESS, one of the only two known multiple planetary system of this kind; Kostov et al. 2020; Standing et al. 2023), virtually “lost” for the time being. For each of all these loose-ephemeris targets, the first two or three months of PLATO photometry will deliver a new, extremely accurate ephemeris, crucial for the community to setup any further follow-up observations.
3.4.2 Transiting candidate planets
Currently, TESS (Ricker et al., 2015) is by far the largest provider of candidate planets within LOPS2. It makes sense to quantify how many. In any case, they will be included into the PIC, not only because some fraction of them [^17] Quantifying the false positive ratio (FPR) of the TOI database is non-trivial because the vetting process is partially done in a manual fashion (Kunimoto et al., 2023). We note, however, that more than 6% of the TOI entries have already been flagged as FP so far, and likely much more will follow before the PLATO launch. will eventually turn out to be genuine planets, but also because PLATO itself (both through its light curves and its follow-up program) can play a fundamental role in their validation and confirmation (Mantovan et al., 2022).
From the 2024-08-20 release of the TESS Object of Interest database (TOI; Guerrero et al. 2021) we identified 824 entries in LOPS2 and matched them with the TESS Follow-up program (TFOP) data base, which includes a disposition flag (“TFOPWG”) reflecting the current status of each target according to the follow-up results. After selecting only successfully vetted planetary candidates (disposition keyword: PC, APC or CP, meaning “planetary candidate”, “ambiguous planetary candidate” and “confirmed planet”, respectively) not already published we are left with 544 candidates hosted by 513 stars [^18] The full list is available on zenodo14720127. (Fig. 7, right panel). We emphasize that these numbers include candidates with a very wide range of FPR, and that a quick inspection at the “comment” column of the TOI database reveals that a considerable fraction of APC entries could likely turn out as false positives. Nevertheless, an individual vetting of each TOI is beyond the scope of this paper. Notably, 27 stars among the LOPS2 TOI host multiple candidates, a configuration which is known to significantly decrease the a-priori FPR. Three systems have multiplicity $N=3$: TOI-699, TOI-790, and TOI-2392 currently under active follow-up by the TFOP working group.
It is worth mentioning that 27 among our list of TESS candidates are listed in the TOI table as having orbital periods longer than 100 d (they are visible at the right end of Fig. 8, right panel) or with no period at all. A closer inspection at the comment field reveals that almost all of them are so called “monotransits”, that is, candidates detected through a single event whose orbital period is unconstrained. In fact, an even larger number of TOI candidates, especially in the long-period end, are flagged as having an ambiguous period due to an insufficient orbital phase coverage by TESS. The almost uninterrupted 2-year coverage of LOPS2 will easily break the ambiguity and detect the correct orbital period (Magliano et al., 2024).
3.4.3 Non-transiting planets
As already mentioned, also non-transiting planets are worth observing with PLATO with the goal of better characterizing the stellar host or to discover additional transiting companions (or both). By cross-matching Exo-MerCat with the sky footprint of LOPS2, after excluding the transiting planets already discussed in the previous sections, we retain a total of 77 planets in 53 systems discovered through three different instrumental techniques (Fig. 9, left plot):
  • Radial velocity: sixty-four planets grouped in 41 planetary systems. Among the 15 multiple systems, HD 40307 (Mayor et al., 2009) stands out as that with the highest multiplicity: at least four planets confirmed by independent analyses, plus two more controversial candidates [^19] As a side note, the inclusion of HD 40307 and other bright RV systems in LOPS2 brings up the opportunity of planning a $\geq 2$ yr observing campaign where ultra-high-precision RV measurements can be combined with a simultaneous and almost uninterrupted space-based light curve to disentangle the planetary signal from stellar activity effects, following the approach developed by Lanza et al. (2011); Aigrain et al. (2012); Haywood et al. (2014), among others. (Tuomi et al., 2013; Díaz et al., 2016). A few RV systems are not counted here because they also host confirmed transiting planets, so they were included in the sample described in Section  3.4.1: they are TOI-500 and TOI-431 (both hosting a transiting super-Earth), L98-59 (with three transiting planets, Kostov et al. 2019; Demangeon et al. 2021), and TOI-1338A = BEBOP-1 (Standing et al., 2023);
  • Direct imaging: eleven planets hosted by ten systems, among which $\beta$ Pic b (Lagrange et al., 2010, 2019), AB Pic (Chauvin et al., 2005), HR 2562 (Konopacky et al., 2016), and WD 0806-661 b (Luhman et al., 2011) stand out as the best characterized. All of them are very massive planets, close to or across the planet-BD transition;
  • Eclipse Timing Variation (ETV): RR Cae ($V=14.4$), a post-common-envelope dM+WD eclipsing binary showing a $P\simeq 16$ years timing modulation claimed as due to the light-travel time effect from an outer $M_{\mathrm{p}}=3.4\pm 0.2$ $M_{\mathrm{jup}}$ giant planet (Qian et al., 2012; Rattanamala et al., 2023). PLATO will obviously be unable to sample the full phase of its signal, yet its uninterrupted series of $\sim 2400$ eclipses will constitute the most precise photometric data set on this target;
  • Astrometry: A sub-stellar candidate orbiting a L1.5 ultracool dwarf, DENIS-P J082303.1-491201 (Sahlmann et al., 2013). At $V\sim 19$ this target is likely too faint to be included in the PLATO target list.
Figure 9: Known non-transiting planetary systems in LOPS2 (orthographic projection in galactic coordinates). Left panel: Entries from the Exo-MerCat database (Section 3.4): planets discovered through RVs (red diamonds), direct imaging (yellow circles), astrometry (gray circle), and ETV (green circle). The location of the candidate TTV system WASP-126 is also marked with a magenta point. Right panel: Candidate astrometric planets published by the Gaia collaboration (purple circles). Each data point is labelled with its ASOI ID number.
It is also worth mentioning WASP-126c, a non-transiting outer companion of the transiting hot Jupiter WASP-126b claimed by Pearson (2019) based on the Transit Timing Variation (TTV) analysis of TESS data, but unconfirmed by subsequent studies (Maciejewski, 2020). If real, its $P\simeq 23$ d, 1-min wide timing modulation would be very easy to confirm or disprove by PLATO, even after a single quarter of photometry.
As anticipated in Paper I, the synergy between PLATO and the thousands of astrometric planets that will be discovered by Gaia (Perryman et al., 2014; Sozzetti et al., 2014) is not to be missed. According to the early estimate by Perryman et al. (2014), 25-50 among those planets, mostly in the short orbital period tail ($P=2$-3  yr), are expected to transit; even for the most favorable cases, the ephemeris will be accurate enough to pinpoint the transit with an error of a few weeks at best (Sozzetti et al., 2023), implying that ground-based campaigns will be largely ineffective or even unfeasible. This opens an exciting opportunity with PLATO, especially for the LOP phase. On top of this, the discovery of additional transiting companions on shorter orbits would be extremely interesting to investigate the architecture of such largely unexplored planetary systems.
The first official release of confirmed astrometric planets from Gaia will be included in DR4, expected not before mid-2026. A preliminary list of candidates has been released in DR3, in the non-single-star (NSS) part (Gaia Collaboration et al., 2023a). Five targets, all M dwarfs, Gaia-ASOI-017, -037, -040, -042, -045 will be monitored in LOPS2 (Fig. 9, right panel). Only Gaia-ASOI-037 ($V\simeq 13.7$), and -040 ($V\simeq 15.4$) are currently included in the PIC; the remaining three are fainter than the $V=16$ requirement set for the PLATO P4 sample. These will possibly be added on a special separate list if simulations will show that meaningful photometry can be extracted.
3.5 Star clusters and associations
It is well known that star clusters provide us with an ideal laboratory to study the formation and evolutionary processes shaping the architecture of planetary systems (Adibekyan et al., 2021). Not only does it provide the opportunity to compare the property of planets hosted in different dynamical environments, also the stellar parameters (including age and chemical composition) of stars belonging to clusters can be measured or derived with a much better accuracy than for field stars (Vejar et al., 2021).
[…]
Figure 10: Open clusters and associations in LOPS2. Left panel: Sky map of the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as red circles with radius r50 (radius containing 50% of the members within the tidal radius). All the 10 682 members at $V<15$ are plotted as black points. Right panel: Eight clusters in our sample having the largest number of $V<15$ targets of spectral type F5V. Each target is color coded with the same scheme as in Fig. 11.
Figure 11: Open clusters and associations in LOPS2. Left panel: the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as a function of their distance and areal density (as defined in the text). The logarithmic age of each cluster is color coded, while the point area is proportional to the number of $V<15$ stars. The eight clusters having the largest number of $V<15$ targets of spectral type F5V and later are labelled. Right panel: Absolute and de-reddened color-magnitude diagram (CMD) for all the 10 6821068210\,68210 682 $V<15$ stars in LOPS2 belonging to clusters (gray points). Members of the eight labeled clusters are plotted with larger circles and color coded as in the legend.
As discussed in Rauer et al. (2024), the main focus of the PLATO planet-hunting survey is on main-sequence stars later than F5V. If we neglect interstellar extinction and assume for the F5V/G2V spectral types an absolute magnitude $M(V)=3.37$ and 4.85, respectively (from Pecaut & Mamajek 2013), we get an upper limit for the distance of 2100 and 1075 pc at a limiting magnitude of $V=15$. Translated at the target level and taking into account the best-fit extinction coefficient AV50 and distance dist50 from Hunt & Reffert (2023), we can set the requirement
$V-\texttt{AV50}+5-5\log(\texttt{dist50})<3.37\texttt{ AND }V<15$ (1)
to get a final list of 3 50635063\,5063 506 targets of interest with spectral type later than F5V. Despite the large number of clusters in LOPS2, more than one third of those targets are hosted by just eight relatively nearby open clusters: NGC 2516, Alessi 34, UPK 545, Haffner 13, NGC 2546, NGC 2451B, IC 2391, NGC 2547, sorted by decreasing number of later than F5V, $V<15$ targets (from 543 to 77) and labelled in the left panel of Fig. 11. They are all located within 1000 pc from the Sun. Their average areal density within r50 (defined as the radius containing half of the cluster members) is much smaller than 1 $\textrm{arcmin}^{-2}$ even for the densest ones, meaning none of them is critically crowded considering the $15^{\prime\prime}$ pixel size of PLATO, with 90% of the flux within $3\times 3$ pixels, and the availability of sophisticated PSF or DIA techniques (Nardiello et al., 2020; Montalto et al., 2020) when a target has been allocated an imagette.
All the mentioned clusters are much younger than the Sun, having estimated ages in the range of 15-150 Myr. This is confirmed by independent spectroscopic studies in the literature, but also by their de-reddened color-magnitude diagrams, showing (with only one exception, see below) a straight main sequence with no hint of evolved members (Fig. 11, right panel). Two clusters in particular stand out because of their properties and deserve some more discussion:
  • IC 2391 is by far the closest object in our sample at a distance of 150 pc. It is a sparse, young ($51\pm 5$ Myr from the Li depletion boundary; Randich et al. 2018) solar-metallicity cluster ($[\mathrm{Fe}/\mathrm{H}]=-0.04\pm 0.03$; De Silva et al. 2013) with about 350 high-confidence members identified so far (Nisak et al., 2022). IC 2391 is also the only cluster in our short list for which late-K stars will be accessible by PLATO. Stellar activity at such young ages could be a limiting factor. For instance, the solar analog Gaia DR3 5318186221414047104 ($V=10.9$, $T_{\mathrm{eff}}=5720$ K, $R_{\star}=0.96$ $R_{\odot}$, $\log g=4.46$), which is a bona-fide member of IC 2391, shows (in TESS light curves) an obvious rotational modulation at $P_{\mathrm{rot}}\simeq 4$ d and with a 3% semi-amplitude. The same star is also a moderately fast rotator, at $v\sin i=9$ km/s (De Silva et al., 2013) requiring an extra effort to get RV confirmation in the rocky planet regime.
  • NGC 2516. The richest cluster of our sample ($\sim$2000 known members, of which 543 at $V<15$ later than F5V; $d=407$ pc), and the only one lying for the most part at $|b|>15^{\circ}$. At an age of $138\pm 40$ Myr (Franciosini et al., 2022) it also is, together with NGC 2546, the oldest cluster of our sample. Metallicity is slightly super-solar, $[\mathrm{Fe}/\mathrm{H}]=+0.08\pm 0.01$ according to Baratella et al. (2020). Interestingly, NGC 2516 has also been proven to possess an extended halo of stars spanning up to 500 pc, or $20^{\circ}$ in the sky (Bouma et al., 2021). As a passing note, we also mention that the hot Jupiter TOI-1937A b (Yee et al., 2023) could possibly belong to this cluster, but the evidence about its membership has been inconclusive so far.
Being a particularly complex task, the selection, characterization and prioritization of the cluster stars to be included in the PIC as a special subset will constitute the main topic of a dedicated future paper.
Other stars in associations, especially young stellar objects (YSO) not yet on the main sequence, could be identified from specific catalogs, such as the Konkoly Optical YSO catalog (KYSO; Marton et al. 2023), a data base of bona-fide, mostly spectroscopically confirmed YSOs compiled by cross-matching Gaia DR3 with several existing catalogs. Among the 11 6711167111\,67111 671 KYSO entries, 431 fall on LOPS2. If we further restrict our sample to $V<15$, only 53 stars survive: 33 belong to the $\gamma$ Velorum cluster (Franciosini et al., 2018), 12 to the Gum Nebula, 5 to different regions/branches of the Vela molecular ridge. Two more stars are field Herbig Ae/Be stars, and one is the brightest star of the open cluster NGC 2362 (Currie et al., 2009), the only one to make it through our magnitude threshold.
3.6 Variable and binary stars
Thanks to its very large area (2149 $\textrm{deg}^{2}$, that is, 5.2% of the whole sky) and to the wide range in galactic latitude covered, LOPS2 will include a large number and variety of variable stars. A cross-match with the latest versions of the VSX (Variable Star indeX; Watson et al. 2006) and Gavras et al. (2023) catalogs yields 47 3564735647\,35647 356 and 282 366282366282\,366282 366 entries, respectively (with notable over-densities in the LMC and close to the galactic plane). The numbers decrease to 31 2113121131\,21131 211 and 15 9321593215\,93215 932, respectively, when we limit our sample at $V<15$. Among these we mention:
  • 7 78777877\,7877 787 eclipsing binaries (including 3017/2155 with explicit detached/contact-type classification);
  • 6 11161116\,1116 111 long-period variables (LPV);
  • 268 $\gamma$ Cassiopeiae stars;
  • 185 chemically peculiar stars;
  • 47 cataclysmic variables (CV);
  • 700 $\delta$ Scuti pulsators;
  • 651 Cepheids and 699 RR Lyr;
  • 29 slowly pulsating B stars (SPB);
  • 17 $\beta$ Cephei stars (BCEP);
  • seven $\gamma$ Doradus stars (GDOR).
It should be emphasized that some of these classes are interesting not just for stellar science, but also to plan more focused planet-hunting surveys. For instance, circumbinary planets represent a rare laboratory to challenge our theories on planetary formation and migration through a disk, and can be discovered both by detecting their transits (Kostov et al., 2020), [^20] The first (and so far only) circumbinary planet detected by TESS, TOI-1338 (Kostov et al., 2020), is by chance in LOPS2. or by modeling their eclipse timing variations (ETV; Goldberg et al. 2023; Brown-Sevilla et al. 2021). Pulsators with intrinsically stable modes, such as $\delta$ Scuti and $\gamma$ Dor variables, may enable the detection of non-transiting planets on wide orbits through the pulsation timing technique, thanks to the unique combination of timing accuracy and temporal baseline of PLATO (Vaulato et al., 2022). Early-type stars with very specific chemical signatures such as $\lambda$ Bootis stars have been predicted since long to host a much higher fraction of giant planets with respect to ordinary field stars, although this has been recently disputed (Saffe et al., 2021).
Some wide-orbit detached eclipsing binaries (DEBs) in LOPS2 can be considered as benchmarks for stellar evolution studies, since their absolute radii and masses can be measured with an extremely high precision ($<3\%$; Serenelli et al. 2021). From the comprehensive list of 273 benchmark DEBs by Southworth (2015), 36 are within LOPS2, and 12 are brighter than $V=15$. Among them CV Velorum ($V=6.69$) a B2.5V+B2.5V system (Albrecht et al., 2014) known to show a significant misalignment between the orbital spin and the rotation axis of its components (Marcussen & Albrecht, 2022), stands out.
Cataclysmic Variables (CVs) are compact binary systems, typically containing a white dwarf which is accreting material from a low-mass star through Roche lobe overflow. Many CVs show outbursts where they brighten by $\sim$2-5 mag on a recurrence time of a few weeks to many months, which is why setting a fixed magnitude limit as done above could be misleading. If we relax that constraint, there are 94 CVs in both VSX and Gavras et al. (2023) which lie in the LOPS2 field, many of which at quiescent are fainter than $V=18$ mag. However, VW Hyi, which shows normal and superoutbursts, and IX Vel which shows Z Cam low states, are relatively bright ($V=12.5$ and $V=10.1$, respectively). Other CVs in the LOPS2 field include UW Pic which is polar, where the white dwarf has a magnetic field strength of $\sim$20 MG. We expect that many CVs and other transients will be readily observable using PLATO during outbursts.
A census of all the SIMBAD objects in LOPS2 sorted by decreasing number of associated publications reveals, besides the objects already mentioned in the previous sections, other specific variable stars of interest:
  • $\gamma$ Doradus ($V\simeq 4.2$), the prototype of the $\gamma$ Dor class of variables (Kaye et al., 1999);
  • AI Velorum ($V\simeq 6.70$), one of the most studied high-amplitude, double-mode $\delta$ Scuti pulsator;
  • AB Doradus, a pre-main-sequence quadruple system (Guirado et al., 2011) known for the super-flaring activity of its primary component (Schmitt et al., 2019), and also the eponymous and brightest member of the AB Dor moving group (Zuckerman et al., 2004);
  • $\gamma^{2}$ Velorum ($V\simeq 1.83$), a binary made of the closest and brightest known Wolf-Rayet (WR; De Marco et al. 2000) star and a blue supergiant, belonging to the Vela OB2 association (Jeffries et al., 2014). Another bright WR star is HR 2583 = HD 50896 ($V=6.91$), a suspected binary system showing brightness variations still of unknown origin (Flores et al., 2023);
  • $\epsilon$ Canis Majoris, a binary with a very bright B2II component ($V\simeq 1.50$), the strongest EUV source in the whole sky;
  • R Doradus, an asymptotic giant branch (AGB) star with a 300 $R_{\odot}$ radius, probably the star with the largest apparent diameter as seen from the Sun (Bedding et al., 1997).
Aside from these few very bright variables and the cross-match with the catalog by Gavras et al. (2023), we have also searched for confirmed non-radial pulsators in LOPS2. This was done from a combined Gaia–TESS approach. Hey & Aerts (2024) distilled about 60 0006000060\,00060 000 variables using light curves from the first year of the TESS mission, starting from the original Gaia DR3 catalog of candidate main-sequence pulsators by Gaia Collaboration et al. (2023b) relying on their stellar properties derived by Aerts et al. (2023). Hey & Aerts (2024) reclassified all the stars whose dominant frequency in the totally independent sparsely sampled Gaia DR3 light curves and the 30-min sampled TESS light curves are equal. All these stars have a dominant amplitude above 4 mmag in the Gaia $G$ band. Among the 6430 stars from Hey & Aerts (2024) occurring in LOPS2, about 5000 are now classified as confirmed multi-periodic pulsators, while the others are rotational variables or eclipsing binaries. These non-radial pulsators in LOPS2 are split up into 1455 gravity-mode pulsators (of $\gamma\,$Dor or SPB type), 2079 $\delta\,$Sct pulsators, and 1449 hybrid pressure- and gravity mode pulsators. Mombarg et al. (2024) provide estimates of the (convective core) masses, radii, and relative ages for more than 10 000 of the $\gamma\,$Dor, SPB, and hybrid pulsators classified by Hey & Aerts (2024). They did so by subjecting the pulsators’ Gaia data to grid-based modelling based on rotating stellar models. These models were constructed such as to be compliant with the asteroseismology of rotating Kepler pulsators in the mass range from 1.3 $M_{\odot}$ and 9 $M_{\odot}$ from Li et al. (2020); Pedersen et al. (2021). Some 3000 of the stars treated by Mombarg et al. (2024) occur in LOPS2. For about 300 among them Aerts et al. (A&A subm.) distilled the near-core rotation frequency from their identified dominant dipole prograde gravito-inertial mode as found consistently in both the Gaia DR3 and TESS light curves. Ongoing work on the full 5-year TESS light curves will increase the number of new pulsators with a measurement of the internal rotation frequency and this estimate will become more precise. Moreover, simulations with PLATOSIM (Jannsen et al., 2024a) reveal that detections of a dense spectrum of oscillation modes suitable for asteroseismology are expected for all these $\gamma\,$Dor and hybrid pulsators brighter than $V<14$ (Jannsen et al., 2024b), making them ideal calibration stars for the stellar science program of the mission.
3.7 Synergy with TESS
Figure 12: Distribution of observed TESS targets in LOPS2, orthographic projection in galactic coordinates. Left panel: All 37 9103791037\,91037 910 CTL targets within LOPS2 observed in short cadence by TESS from sector 1 to 82 included (black points), superimposed on the LOPS2 field. The yellow circle is the TESS southern CVZ (up to 35 sectors), the red circle is roughly the northern boundary of Camera 4, that is, the sky region where at least 4 TESS sectors are available. Right panel: Same but the TESS FFI coverage in terms of TESS sectors is color coded over an HEALPix level-6 grid.
It is worth investigating the synergy with TESS not only in terms of known planets and planetary candidates, but also because the availability of long ($\geq$ 28 d) and precise light curves can enable a useful characterization of the target stars in advance, including parameters which can play a role at the prioritization stage, such as rotational periods, activity levels, variability and/or binarity.
Among the 519 704519704519\,704519 704 unique TESS targets (Stassun et al., 2018, 2019) observed at the regular 2-min cadence up to Sector 82 included (i.e., August 2024, end of Cycle 6), 37 9103791037\,91037 910 lie within the approximate footprint of LOPS2 with an average areal density spanning the 5-20 stars per $\textrm{deg}^{2}$ range (Fig. 12, left panel; the two most evident over-densities are centered on the LMC and NGC 2516). Among these, 55% were observed for at least three TESS sectors (not always contiguous) and 15% for at least 13 sectors, or one year cumulated (mostly within the $\beta\lesssim-78^{\circ}$ cap). Slightly more than 40% of the available TESS short-cadence targets are not included in the current version of the PIC, [^21] Being included in the PIC does not automatically imply being scheduled for observations or being included in the P1-P2-P4-P5 sample, even though the vast majority of them will do (Montalto et al., 2021). mostly because their stellar parameters fall outside our $(T_{\mathrm{eff}},R_{\star})$ parameter space (M21), or, to a lesser extent, because they do not meet our magnitude requirements. All those remaining targets of scientific interest left out can be requested by the community through the PLATO GO program, if they are technically feasible.
If we focus on the general sector coverage of the TESS FFIs instead (Fig. 12, right panel) we find that 25% of the LOPS2 sky area has been already surveyed through at least 12 TESS sectors, 55% through at least five sectors, and 88% by three.
3.8 Synergy with CHEOPS
CHEOPS (Benz et al., 2021) is an ESA S-class mission launched in 2019, equipped with a 30-cm reflecting telescope designed for ultra-high-precision photometry of individual targets in a single optical band (Fortier et al., 2024). Its main goal is the follow-up and characterization of known transiting planets, articulated over a wide range of science cases; see Benz et al. (2021) for a summary of the CHEOPS GTO program. Its 3.5-year primary mission ended in 2023, and currently CHEOPS is running its first mission extension until 2026, with a possible second extension foreseen until 2029. Given the scientific (and possibly temporal) overlap with the PLATO observations, it is worth investigating how much of LOPS2 can be accessed by CHEOPS.
CHEOPS is not subject to a rigid scanning law such as TESS. Rather, it works on a flexible schedule, and its pointing ability is limited only by three avoidance angles to minimize scattered light from the Sun, the Moon and the Earth limb. The Sun Exclusion Angle (SEA) is currently set to $120^{\circ}$ (as throughout the whole nominal mission), and is the most limiting factor to reach large ecliptic latitudes such as those spanned by LOPS2. It is easy to see that a SEA of 120 implies that the whole $\beta<-60^{\circ}$ cap (the magenta area plotted in Fig. 13, left panel) is inaccessible to CHEOPS. In other words, only 33% of the LOPS2 footprint can currently be pointed by CHEOPS, mostly in the six- and 12-camera regions and in any case for a very limited amount of days per year, because of the SEA limitation. The 24-NCAM region is completely inaccessible. Moreover, the fraction of accessible P1 targets (black dots in Fig. 13) is even lower (23%), since P1 targets are more densely located in the inner regions (18 and 24 telescope) of LOPS. Should the CHEOPS observing strategy and/or the value of the avoidance angles be changed during the extended mission, the geometric overlap could increase by a significant amount; a relaxation of the SEA of a few degrees, for instance, would make part of the PLATO 24-camera region and its dense population of P1 targets accessible to CHEOPS.
Figure 13: Overlap between CHEOPS/Ariel and LOPS2 (orthographic projection in galactic coordinates). Left panel: Sky area forbidden to CHEOPS (in magenta) due to the Sun Exclusion Angle currently set at $120^{\circ}$ (corresponding to $\beta<-60^{\circ}$ for the southern ecliptic cap). P1 targets are plotted as black dots. Right panel: Ariel targets from the current MRS. Confirmed targets belonging to the Ariel Tier 1/2/3 (see text for details) are plotted as yellow, orange and red dots, respectively, while the TESS candidate planets are plotted as gray crosses (see text for details). The yellow circle is the TESS CVZ.
3.9 Synergy with Ariel
Ariel is an ESA M-class mission currently in development, with a planned launch to L2 in 2029 (Tinetti et al., 2018, 2021), i. e., with a significant temporal overlap with the PLATO nominal mission. Based on a 1-m class telescope equipped with two near-to-mid infrared low-resolution spectrographs (covering the 1.1-7.8 $\mu$m range) and three photometric channels also employed as fine guidance system (FGS; 0.5-1.1 $\mu$m), Ariel will monitor transiting planets through emission spectroscopy, transmission spectroscopy and phase curves, with repeated observations to gather the required SNR. Ariel targets will be observed through a four-“Tier” approach: Tier 1 targets will undergo an extensive low-resolution reconnaissance survey, Tier 2 targets will be selected for a deeper survey at higher resolution, and Tier 3 is made of a relatively small number of benchmark planets to be explored for variability (Edwards & Tinetti, 2022). Tier 4 will be dedicated to phase curves and bespoke observations of targets of special interest, which are expected to be identified, for example, in Tier 1 or by space-based missions and ground-based surveys prior to Ariel.
The current Ariel target list, called Mission Reference Sample (MRS) is a living catalog, started by Edwards & Tinetti (2022) and regularly updated. [^22] https://github.com/arielmission-space/Mission_Candidate_Sample The latest version (2024-07-09) contains 722 confirmed transiting planets, and 2025 candidates from the TESS TOI/ExoFOP database (Guerrero et al., 2021). The majority of MRS targets (58%) are hot Jupiters. Also, nearly 90% of the “confirmed planets” sub-sample have transits deeper than 1 mmag, i. e., are easily detectable by TESS at high SNR (or in most cases, even from ground-based facilities). A cross-match with the LOPS2 footprint reveals that PLATO will observe 60 confirmed planets from the MRS (22 in Tier-1, 29 in Tier-2, nine in Tier-3) belonging to 49 host stars (Fig. 13, right panel).
Given the selection criteria of the MRS, PLATO will be a provider of new, additional targets for Ariel, mainly of a small (but scientifically very interesting) population of long-period transiting giants not (yet) discovered by TESS, or discovered as mono-transit events only (Magliano et al., 2024). Moreover, the PLATO versus Ariel synergy will be strong, for the already known and future targets, on the extremely accurate stellar parameters delivered by PLATO through asteroseismological analysis (and in particular, ages; Goupil et al. 2024), the in-depth characterization of stellar activity and rotational periods (Breton et al., 2024), the refinement of the planetary ephemeris for targets difficult to be followed up from the ground, the extension of the temporal baseline for TTV studies (Borsato et al., 2022), and the accurate planetary masses obtained by the PLATO follow-up program (Rauer et al., 2024). On the hot Jupiters belonging to the brighter tail of the MRS, PLATO will also be able to detect the planetary albedo and/or heat redistribution behavior through phase curves (Shporer, 2017; Singh et al., 2019), in a few cases possibly with color information thanks to the FCAM coverage.
3.10 Synergy with JWST
On a closing note, we also mention that the JWST southern CVZ at $\beta<-85^{\circ}$ (Gardner et al., 2006) is fully enclosed in the LOPS2 footprint, unfolding another synergy with a space-based mission that is and will be crucial in the investigation of exoplanetary atmospheres. Further, JWST is able to monitor continuously for at least 200 days every southern source at $\beta<-45^{\circ}$, and this includes the whole LOPS2.
4 Conclusions
In this work we presented the first field to be observed by PLATO during its LOP phase, and illustrated some of its astrophysical content relevant for planetary and stellar science. While the position of LOPS2 on the sky is now fixed, the PIC will continue to evolve and improve. The release of Gaia DR4 will deliver much more accurate stellar parameters for all of our targets of interest, including all non-single star (NSS) solutions, individual epoch measurements, variability metrics, metallicities/abundances etc. based on the first 66 months of observations, and a large catalog of planets and planetary candidates discovered by Gaia astrometry. As presented in this paper, LOPS2 also contains numerous highly relevant targets to address complementary science topics. In this context, the general community will be invited to propose observations in response to ESA announcements of opportunity for a Guest Observer program. The first open call is planned to be issued nine months prior to launch.
Meanwhile, effort will be spent on building the more effective target lists and defining criteria on how to allocate the PLATO telemetry resources (imagettes, centroids, light curves) among the selected targets. The selected field meets the mission scientific requirements on samples P1-P2-P4-P5 as defined by SciRD. In the near future, the PSWT will implement avprioritization metric able to define a subset of most valuable stars for the ground-based follow-up, called prime sample. For planets hosted by this sample, the PLATO mission will eventually provide planet candidate confirmation and high-resolution spectroscopy, with planetary mass measurements. The prioritization process will also allow prioritization of tPIC targets for imagettes acquisition, sampling timing, etc. This process, to be documented in a third paper of this series, does not separately consider the different samples. The prioritization metric is evaluated regardless of whether a star belongs to the P1, P2 or P5 sample. For instance, a K3 subgiant at $V\lesssim 11$ and $R_{\star}\simeq 3$ $R_{\odot}$ can be included in P1 with a similar $\textrm{NSR}\simeq 45$ ppm with respect to a K3V star at $V\gtrsim 11$ and $R_{\star}\simeq 0.8$ $R_{\odot}$ belonging to P5, due to the magnitude constraint. However, detecting a true Earth analog during LOPS would be impossible in the former case, while perfectly within the reach [^23] As a figure of reference, a photometric precision of 80 ppm in one hour enables the robust detection of a true Earth analog around a quiet Sun twin with three transits (Rauer et al., 2024). of PLATO in the latter, so that this observation will be given higher priority.
In synthesis, the prioritization scheme will lead to the selection of the PLATO prime sample, a PIC subset of up to 20 0002000020\,00020 000 stars to be observed during the LOP for which the PLATO Consortium is committed to do and deliver the follow-up observations. For that purpose, additional constraints (also to be documented in Paper III) other than the prioritization metric will be added to ensure that the targets are feasible with the available ground-based facilities and within the timeline of the PLATO operations.
Data availability
The table with the confirmed transiting planets in LOPS2, the table with the TESS candidates in LOPS2 and the MOC regions of LOPS2 are publicly available on Zenodo: zenodo14720127.
Acknowledgements.
We thank the referee for his valuable comments and suggestions. We thank Hans Deeg for his useful comments.
This work presents results from the European Space Agency (ESA) space mission PLATO. The PLATO payload, the PLATO Ground Segment and PLATO data processing are joint developments of ESA and the PLATO mission consortium (PMC). Funding for the PMC is provided at national levels, in particular by countries participating in the PLATO Multilateral Agreement (Austria, Belgium, Czech Republic, Denmark, France, Germany, Italy, Netherlands, Portugal, Spain, Sweden, Switzerland, Norway, and United Kingdom) and institutions from Brazil. Members of the PLATO Consortium can be found at https://platomission.com/. The ESA PLATO mission website is https://www.cosmos.esa.int/plato. We thank the teams working for PLATO for all their work.
CA and AT acknowledge support from the long-term structural Methusalem funding program by means of the project SOUL: Stellar evolution in full glory, KU Leuven grant METH/24/012. JMMH is funded by Spanish MCIU/AEI/10.13039/501100011033 grant PID2019-107061GB-C61. VN, GP, MM, SB, SD, VG, DM, LM, IP, LP, RR acknowledge support from PLATO ASI-INAF agreements n. 2022-28-HH.0
This research has made use of the SIMBAD database (operated at CDS, Strasbourg, France; Wenger et al. 2000), TOPCAT and STILTS (Taylor, 2005, 2006), UCAC4-RPM (Nascimbeni et al., 2016), TEPCAT (Southworth, 2011). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI:10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
Appendix A The LOPS2 footprint
A software tool to check whether a given line of sight passes through the PLATO fields will be made available in the near future by ESA with support from the PLATO Consortium. Meanwhile, for any non-critical purpose, the footprint of LOPS2 can be reasonably approximated by the intersection between a spherical circle with $r=28.1^{\circ}$ and a “square” made with arcs of great circles (cf. Fig. 1). The corresponding TOPCAT expression is:
inSkyEllipse(l,b,255.9375,-24.62432,28.1,
28.1,0) && inSkyPolygon(l,b,288.6075,-44.1, (2)
223.2675,-44.1,233.1675,-0.25,278.7075,-0.25)
where $(l,b)$ are the galactic coordinates of the target. This expression can be translated into any other language of choice since it contains the coordinates of the five pivot points involved (the center of the circles plus the four vertexes of the square).
A more accurate representation of LOPS2, also including the NCAM sub-regions, can be built by approximating the field on a level-9 HEALPix grid, and then converting the footprints into IVOA [^24] International Virtual Observatory Alliance, https://ivoa.net/-compliant Multi-Order Coverage (MOC) regions (Fernique et al. 2014), through the mocpy code [^25] https://cds-astro.github.io/mocpy/. The MOC regions for 6, 12, 18, 24 NCAMs and the underlying HEALPix grid are available online262626zenodo14720127. We emphasize that the actual sky coverage could be slightly different due to several factors (the in-orbit optical performances, the relative co-alignment of the cameras, the pointing accuracy, etc.) so all these regions should be used with some caution especially when a target of interest is close the the external boundaries or to the inner gaps.
Appendix B Additional tables and figures
Table 3: Glossary of acronyms used throughout this article.
Acronym Description
asPIC All-sky PLATO Input Catalog (M21)
CVZ Continuous Viewing Zone
DEB Detached Eclipsing Binary
DIA Difference Image Analysis
EB Eclipsing Binary
FPR False Positive Rate
FCAM PLATO Fast camera
FOV Field Of View
GC Globular Cluster
GO PLATO Guest Observing program
HJ hot Jupiter
HZ Habitable Zone
LMC Large Magellanic Cloud
LOP Long-duration Observation Phase
LOPN LOP field North
LOPN1 Current LOPN proposal (this work)
LOPS LOP field South
LOPS1 Current LOPS proposal (this work)
M21 Montalto et al. (2021)
N22 Nascimbeni et al. (2022)
NCAM PLATO Normal camera
NSR Noise-to-Signal ratio
ORM Observatorio Roque de Los Muchachos
OC Open Cluster
P/L PLATO Payload
PF PLATO field
PIC PLATO Input Catalog (M21)
PLATO PLAnetary Transits and Oscillations of stars (Rauer et al. 2024)
PPT PLATO Performance Team
PSF Point Spread Function
RV Radial velocity
SNR Signal-to-Noise ratio
SOP Step and stare Observation Phase
SciRD PLATO Scientific Requirements Document
SRJD PLATO Scientific Requirements Justification Document
SWT ESA PLATO Science Working Team
TCP TESS candidate planets
TTV Transit timing variations
WJ warm Jupiter
Figure 14: Optimization of the PLATO field rotation angle. The number of P1 (left panel) and P5 (right panel) targets are shown as a function of the rotation angle, for hypothetical PLATO fields centered at the LOPN1 (blue points and lines) or LOPS2 (green points and lines) coordinates. The continuous and dashed horizontal lines mark, for reference, the value of P1/P5 counts at $\varphi=0^{\circ}$ and a ±$\pm 1\%$ deviation from it, respectively (see Section 2.2 for more details).
Figure 15: Optimization of the PLATO field rotation angle. The PLATO test fields are centered at the LOPS2 coordinates (Table 2) but rotated by $0^{\circ}$, $10^{\circ}$, $20^{\circ}$, $30^{\circ}$, $40^{\circ}$, $50^{\circ}$, $60^{\circ}$, $70^{\circ}$, $80^{\circ}$ (in reading order); projection is galactic. The $\varphi=0^{\circ}$ rotation angle corresponds to the nominal LOPS2 geometry. As in Fig. 6, the brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3). The TESS southern CVZ is marked with a yellow circle (see Section 2.2 for more details).
[…]
Figure 17: All-sky Aitoff projection in Equatorial coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
Figure 18: All-sky Aitoff projection in Ecliptic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
Figure 19: Distribution in distance and interstellar extinction of LOPS2 targets, from the PIC. Both the distribution of the whole sample (in gray) and for a few subsamples of interest is shown: M dwarfs, i. e., P4 (in red), main-sequence F stars (blue), main-sequence G and K stars (green) and FGK subgiants (orange). The definition of these sub-samples in terms of $T_{\mathrm{eff}}$ and $R_{\star}$ can be found in N22, Section 6.3. Left panel: Logarithmic histogram of distance. Right panel: Logarithmic histogram of the interstellar extinction in the Bessel $V$ band.
Figure 20: Distribution in $V$, Gaia $G$, 2MASS $J$, 2MASS $K_{s}$ magnitude (panels from upper left to lower right) of LOPS2 targets, from the PIC. Both the distribution of the whole sample (in gray) and for a few subsamples of interest is shown: M dwarfs, i. e., P4 (in red), main-sequence F stars (blue), main-sequence G and K stars (green) and FGK subgiants (orange).
Figure 21: Histograms of some astrophysical quantities of the P1+P2 (red histograms, left side) and P5 (blue histograms, right side) samples within the LOPS2 field (see Section 3.1 for details). From top to bottom: effective temperature $T_{\mathrm{eff}}$ in K, stellar radius $R_{\star}$ and mass $M_{\star}$ in solar units, and orbital period $P_{\mathrm{HZ}}$ corresponding to the center of the habitable zone according to the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=\sqrt{L_{\star}/L_{\odot}}$) and the Kepler’s laws. The Sun/Earth values are marked with dashed vertical line as reference.
Figure 22: Orbital periods corresponding to the habitable zone around PLATO targets. Upper panels: Histograms of the expected orbital period of a planet at the center of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=\sqrt{L_{\star}/L_{\odot}}$) and Kepler’s laws, for the P1+P2 (left side) and P5 (right side) samples within the LOPS2 field (see Section 3.1 for details). The Sun/Earth value (1 yr) is marked with dashed vertical line as reference. Both the distribution of the whole sample (in gray) and for a few subsamples of interest is shown: main-sequence F stars (blue), main-sequence G and K stars (green) and FGK subgiants (orange). The definition of these sub-samples in terms of $T_{\mathrm{eff}}$ and $R_{\star}$ can be found in N22, Section 6.3. Lower panels: Same but for a planet located at the inner edge of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=0.77\times\sqrt{L_{\star}/L_{\odot}}$).
Figure 23: Spatial distribution of the P1-P2-P4-P5 targets (in reading order) over the LOPS2 field. P1, P4 and P5 targets are represented as color coded density maps based on an HEALPix level-5, while the sparse and almost perfectly isotropic distribution of the P2 sample is plotted with individual red points. See also Section 3.1 for a discussion of the sample properties.
1]
Table 4: List of confirmed transiting exoplanets in LOPS2 (see Section 3.4.1 for details).
planet name $N_{p}$ $P$ [days] $R_{\mathrm{p}}/R_{\oplus}$ $M_{\mathrm{p}}/M_{\oplus}$ $T_{\mathrm{eq}}$ [K] $V$ [mag] edge? $N_{\mathrm{cam}}$ FCAM? comment
GJ 238 b 1 1.74 0.6 750 11.57 no 24 yes
GJ 341 b 1 7.58 0.9 532 9.46 yes 6 no
HATS-39 b 1 4.58 17.6 200.2 1635 12.74 no 6 no
HATS-40 b 1 3.26 17.7 505.3 2085 13.48 no 12 no
HATS-41 b 1 4.19 14.9 3083.0 1683 12.68 no 6 no
HATS-42 b 1 2.29 15.7 597.5 1842 13.68 no 12 no
HATS-43 b 1 4.39 13.2 83.0 993 13.56 no 12 no
HATS-44 b 1 2.74 12.0 178.0 1164 14.4 no 12 no
HATS-45 b 1 4.19 14.4 222.5 1508 13.32 no 6 no
HATS-51 b 1 3.35 15.8 244.1 1540 12.52 no 6 no
HATS-55 b 1 4.2 14.0 292.7 1359 13.52 no 6 no
HATS-66 b 1 3.14 15.8 1694.0 1984 14.28 no 12 yes
HATS-70 b 1 1.89 15.5 4100.0 2723 12.23 no 6 no UHJ
HATS-76 b 1 1.94 12.1 835.6 935 16.68 no 6 no very faint
HD 21749 b 2 35.61 2.6 22.7 421 8.08 yes 6 no =GJ143 b
HD 21749 c 2 7.79 0.9 3.7 698 8.08 yes 6 no =GJ143 c
HD 23472 b 5 17.67 2.0 8.3 547 9.73 no 6 no
HD 23472 c 5 29.8 1.9 3.4 465 9.73 no 6 no
HD 23472 d 5 3.98 0.8 0.6 908 9.73 no 6 no
HD 23472 e 5 7.91 0.8 0.7 722 9.73 no 6 no
HD 23472 f 5 12.16 1.1 0.8 627 9.73 no 6 no
HD 28109 b 3 22.89 2.2 18.5 917 9.42 no 6 no
HD 28109 c 3 56.01 4.2 7.9 609 9.42 no 6 no
HD 28109 d 3 84.26 3.2 5.7 527 9.42 no 6 no
HD 56414 b 1 29.05 3.7 1126 9.22 no 12 no
KELT-14 b 1 1.71 19.5 408.1 1953 11.0 no 24 no
KELT-15 b 1 3.33 19.5 416.4 1775 11.39 no 12 no
KELT-25 b 1 4.4 18.4 2265 9.84 no 6 no UHJ
L 98-59 b 3 2.25 0.8 0.4 614 11.68 yes 6 no
L 98-59 c 3 3.69 1.4 2.2 532 11.68 yes 6 no
L 98-59 d 3 7.45 1.5 1.9 410 11.68 yes 6 no
LHS 1678 b 3 0.86 0.7 857 12.6 no 6 no USP
LHS 1678 c 3 3.69 0.9 528 12.6 no 6 no
LHS 1678 d 3 4.97 1.0 477 12.6 no 6 no
LHS 1815 b 1 3.81 1.1 1.6 610 12.17 no 12 yes
NGTS-1 b 1 2.65 14.9 258.1 785 15.67 no 12 no very faint
NGTS-10 b 1 0.77 13.5 687.1 1432 14.51 no 12 no USP
NGTS-15 b 1 3.28 12.3 238.7 1192 14.67 no 6 no
NGTS-17 b 1 3.24 13.9 242.8 1516 14.41 no 6 no
NGTS-23 b 1 4.08 14.2 194.8 1379 14.13 no 12 no
NGTS-29 b 1 69.34 9.6 124.9 439 10.51 no 6 no
NGTS-3 A b 1 1.68 16.6 756.4 1695 14.67 no 12 yes
NGTS-4 b 1 1.34 3.2 20.6 1639 13.14 no 12 no
NGTS-6 b 1 0.88 14.9 425.6 1395 14.24 no 6 no USP
TOI-1221 b 1 91.68 2.9 1112.4 437 10.49 no 12 no
TOI-1338 b 1 95.4 7.7 11.3 487 11.72 no 18 yes
TOI-163 b 1 4.23 16.7 387.8 1659 11.47 no 6 no
TOI-1937 A b 1 0.95 14.0 638.8 2083 13.18 no 18 no USP
TOI-199 b 1 104.85 9.1 54.0 347 10.7 no 18 no TTV
TOI-201 b 1 52.98 11.3 133.5 699 9.07 no 24 yes
TOI-206 b 1 0.74 1.3 905 14.94 no 6 no USP
TOI-216.01 2 34.53 10.1 178.0 476 12.32 no 6 no
TOI-216.02 2 17.16 8.0 18.8 603 12.32 no 6 no
TOI-2184 b 1 6.91 11.4 206.6 1429 12.25 no 12 no
TOI-220 b 1 10.7 3.0 13.8 794 10.47 no 12 yes
TOI-2338 b 1 22.65 11.2 1900.6 742 12.48 no 12 no
TOI-2368 b 1 5.18 10.8 206.6 995 12.49 no 6 no
TOI-2416 b 1 8.28 9.9 953.5 1074 13.02 no 6 no
TOI-2459 b 1 19.1 3.0 445 10.77 no 18 no
TOI-2525 b 2 23.29 8.7 26.7 555 14.22 no 12 yes
TOI-2525 c 2 49.25 10.1 208.8 432 14.22 no 12 yes
TOI-2529 b 1 64.59 11.6 743.0 633 11.53 no 12 no
TOI-2589 b 1 61.63 12.1 1112.4 547 11.42 no 12 no
TOI-269 b 1 3.7 2.8 8.8 550 14.37 no 12 no
TOI-270 b 3 3.36 1.2 1.6 578 12.6 no 12 no
TOI-270 c 3 5.66 2.4 6.2 486 12.6 no 12 no
TOI-270 d 3 11.38 2.1 4.8 385 12.6 no 12 no
TOI-2803 A b 1 1.96 18.1 309.9 1883 12.54 yes 6 no
TOI-2818 b 1 4.04 15.3 225.7 1368 11.94 no 12 no
TOI-286 b 2 4.45 1.4 4.6 972 9.87 no 12 yes
TOI-286 c 2 39.39 1.9 3.7 472 9.87 no 12 yes
TOI-431 b 2 0.49 1.3 3.1 1859 9.12 no 6 no USP
TOI-431 d 2 12.46 3.3 9.9 632 9.12 no 6 no
TOI-451 b 3 1.86 1.9 1482 10.94 no 6 no
TOI-451 c 3 9.19 3.1 869 10.94 no 6 no
TOI-451 d 3 16.36 4.1 718 10.94 no 6 no
TOI-4562 b 1 225.12 12.5 732.0 337 12.14 no 12 no very long P
TOI-470 b 1 12.19 4.3 656 11.17 no 12 no
TOI-481 b 1 10.33 11.1 486.3 1250 9.97 no 12 no
TOI-500 b 1 0.55 1.2 1.4 1608 10.54 no 24 yes USP
TOI-540 b 1 1.24 0.9 607 14.82 no 12 yes
TOI-622 b 1 6.4 9.2 96.3 1376 9.0 no 12 no
TOI-640 b 1 5.0 19.9 279.7 1737 10.51 no 18 yes
TOI-700 b 4 9.98 0.9 410 13.15 no 12 yes
TOI-700 c 4 16.05 2.6 350 13.15 no 12 yes
TOI-700 d 4 37.42 1.1 264 13.15 no 12 yes HZ
TOI-700 e 4 27.81 1.0 292 13.15 no 12 yes HZ
TOI-712 b 3 9.53 2.0 645 10.84 no 12 yes
TOI-712 c 3 51.7 2.7 367 10.84 no 12 yes
TOI-712 d 3 84.84 2.5 311 10.84 no 12 yes HZ
TOI-813 b 1 83.89 6.7 606 10.36 no 6 no
TOI-858 B b 1 3.28 14.1 349.6 1520 11.18 yes 6 no
TOI-871 b 1 14.36 1.7 617 10.57 no 12 no
WASP-100 b 1 2.85 14.9 400.5 1985 10.8 no 6 no
WASP-101 b 1 3.59 16.0 162.1 1545 10.34 no 6 no
WASP-119 b 1 2.5 15.7 390.9 1573 12.31 no 6 no
WASP-120 b 1 3.61 16.5 1541.5 1866 10.96 no 12 no
WASP-121 b 1 1.27 19.6 367.7 2306 10.51 no 12 no UHJ
WASP-126 b 1 3.29 10.8 90.3 1468 10.99 no 6 no TTV claim
WASP-159 b 1 3.84 15.5 174.8 1837 12.84 no 6 no
WASP-160 B b 1 3.77 12.2 88.4 1113 13.04 no 12 no
WASP-168 b 1 4.15 16.8 133.5 1333 12.12 no 24 yes
WASP-23 b 1 2.94 10.8 281.0 1109 12.54 no 24 yes
WASP-61 b 1 3.86 15.8 851.8 1623 12.49 no 6 no
WASP-62 b 1 4.41 14.8 165.3 1381 10.21 no 12 yes
WASP-63 b 1 4.38 15.8 117.6 1504 11.16 no 24 yes
WASP-64 b 1 1.57 14.2 403.9 1658 12.7 no 12 yes
WASP-79 b 1 3.66 17.2 270.2 1705 10.04 yes 6 no
Table 4: Continued.
 [^27] The columns give: the name of the planet, the multiplicity of the system, the orbital period $P$ in days, the planetary radius $R_{\mathrm{p}}$ and mass $M_{\mathrm{p}}$ in Earth units (when known), the planetary equilibrium temperature in K, the $V$ magnitude of the host star, a flag if the target is close to the LOPS2 outer edge, the number of NCAMs, a flag if the target is also observed with the FCAMs, and a comment field.
2]
Table 5: List of TESS candidates in LOPS2, from the TOI/ExoFOP database (see Section 3.4.2 for details).
TOI ID $N_{p}$ $T$ [mag] $P$ [days] $R_{\mathrm{p}}/R_{\oplus}$ $\delta$ [ppm] $T_{\mathrm{eq}}$ [K] comment
119.01 2 9.23 5.541 2.04 633 1117 Potential multi; slight depth aperture correlation with no offset
119.02 2 9.23 10.692 1.8 535 710 outer candidate in potential multi; variable host
124.01 1 11.04 1.843 21.27 1950 1785 TFOP SB1/APC; v-shaped
153.01 1 11.69 7.632 12.44 4580 0
163.01 1 10.87 4.231 14.63 7261 1584 TOI-163 b
167.01 1 11.57 4.453 0.0 5910 1027 no stellar radius; Gaia DR2 Teff + plx indicate sun-like star
171.01 1 11.54 1.239 13.58 13300 1461
174.01 5 8.7 17.667 1.87 854 462 HD 23472 b / TOI 174.01
174.02 5 8.7 29.798 1.61 671 374 HD 23472 c; multi
174.03 5 8.7 12.162 1.16 317 638 HD 23472 f / TOI 174.03
174.04 5 8.7 3.977 0.76 142 855 HD 23472 d / TOI 174.04
174.05 5 8.7 7.908 0.97 217 720 HD 23472 e/ TOI 174.05
187.01 1 9.09 0.513 19.23 5691 3118 v-shaped; very likely EB; TFOP APC previously retired as BEB due to chromaticity; possible odd-even
189.01 1 10.22 2.194 8.01 5980 1332 TFOP APC/BEB; v-shaped
199.01 1 10.0 104.873 10.22 12183 358 TOI-199 b (https://arxiv.org/pdf/2309.14915.pdf)
201.01 2 8.57 52.978 11.45 6347 598 TOI 201 b
201.02 2 8.57 5.849 1.19 109 1237 TOI 201.02; interior to TOI-201 b
206.01 1 12.47 0.736 1.53 1323 840 TOI-206 b
211.01 1 9.45 6.564 7.63 2132 1003 TFOPWG APC/BEB
214.01 2 8.01 18.554 1.58 448 623 L1 candidate; possible multi
214.02 2 8.01 9.696 0.81 122 794 inner candidate in possible multi
216.01 2 11.5 34.506 10.46 17388 449 TOI-216.01/TOI-216 c; clear TTVs in spoc-s1-s39 multisector
216.02 2 11.5 17.389 6.48 2972 564 TOI-216.02 / TOI-216b. Apparent TTVs. Possible additional single around 2046P
218.01 2 13.55 0.438 0.65 486 901 Bad transit shape; likely SV
218.02 2 13.55 8.352 1.8 3925 337 eruptive variable host; TOI 218.01 (Psimilar-$\sim$0.48 d) may be SV
220.01 1 9.66 10.695 2.85 920 796 TOI 220 b (Hoyer et al 2021)
252.01 1 11.62 1.002 5.02 2320 1563 TFOP APC; synchronized to variability
268.01 1 9.6 5.066 7.69 1350 1071 TICID 219253008 is a phantom from split (TICID 685340264 and 685340263)
269.01 1 12.27 3.698 2.94 4756 534 TOI-269 b
270.01 3 10.42 5.661 2.41 3856 453 TOI-270 c / TOI 270.01
270.02 3 10.42 11.38 2.25 3147 359 TOI-270 d / TOI-270.02
270.03 3 10.42 3.36 1.18 973 539 TOI-270 b / TOI 270.03
271.01 1 8.41 2.476 2.81 387 1606 follow up in progress
275.01 1 10.98 0.92 8.7 9180 1913 TFOP APC/BEB; v-shaped
281.01 1 10.48 5.577 4.56 642 1293 L1 candidate
282.01 3 8.93 56.005 4.61 768 597 HD 28109 c
282.03 3 8.93 84.261 3.44 440 521 HD 28109 d / TOI 282.03
282.04 3 8.93 22.891 2.56 253 804 HD 28109 b; multi
283.01 1 9.65 17.617 2.01 722 610
285.01 1 12.11 32.333 3.03 2426 329 possible single transit
286.01 2 9.08 4.512 1.4 313 850 potential L1 planet; potential multi
286.02 2 9.08 39.362 1.85 552 606 potential multi; L1 candidate
322.01 1 10.71 3.991 16.08 3306 1382
386.01 1 10.01 5.112 10.67 1415 1521 v-shaped; two stars about 1” separation in same pixel; slight depth aperture correlation
407.01 1 12.7 2.165 13.56 4990 1970 TFOP SB1/APC
429.01 1 10.17 0.0 9.33 7526 958 single transit with unknown period; still a single transit as of Sector 69
431.01 2 8.17 12.461 3.28 2078 589 TOI-431 d / TOI 431.01
431.02 2 8.17 0.49 1.47 291 1731 TOI-431 b / TOI 431.02
445.01 1 9.14 0.765 12.95 1900 2971
446.01 1 11.88 3.502 22.1 16196 1648 TFOP APC; retired as SB1
447.01 1 8.88 5.529 23.55 17914 1235 big;
448.01 1 13.17 0.882 8.39 11755 1428 NGTS-6 b
450.01 1 12.38 10.715 14.04 56326 340 TFOP APC; retired as SB2
451.01 3 10.27 16.365 3.92 2075 634 TOI-451 d / TOI-451.01
451.02 3 10.27 1.859 1.93 498 1310 TOI-451 b
451.03 3 10.27 9.191 2.73 941 769 TOI-451 c
459.01 1 9.54 4.429 6.34 700 1728 TFOP APC/SB1; synchronized to variability
470.01 1 10.65 12.191 4.61 2625 660 TOI-470 b
472.01 1 12.29 0.568 24.85 42540 954 no GAIA radius with Teff 3869K; large planet around an M star if real
475.01 1 10.77 8.262 2.34 1418 531
476.01 1 8.85 3.063 14.75 6160 1930 pulsating host
481.01 1 9.37 10.331 11.94 4577 $1$ TOI-481 b
486.01 1 9.34 1.745 0.59 160 896 high proper motion star
499.01 1 10.1 8.533 4.15 1756 941 v-shaped; possible secondary; possible centroid offset towards TIC 769340768
500.01 1 9.32 0.548 1.36 227 1577 TOI-500 b
502.01 1 10.33 2.941 8.83 2421 1431 odd-even; V-shaped; possible EB
510.01 1 8.44 1.352 2.83 371 1110 TFOP RV0/APC; variable host; offset not consistent with nearby stars
512.01 2 8.95 7.189 1.53 285 831 first candidate in potential multi
512.02 2 8.95 20.275 1.74 410 764 potential multi
520.01 1 9.37 0.524 1.49 83 3256
527.01 1 8.24 18.089 22.2 5527 984 planet radius greater than 20 Re which is large for orbital period; v-shaped
533.01 1 10.89 19.572 3.61 1677 540
539.01 1 10.54 0.31 1.55 258 2138 a/Rssimilar-$\sim$2; possibly evaporating
540.01 1 11.48 1.239 0.97 2319 550 TOI-540 b
541.01 1 14.6 1.353 18.11 18071 1618 low SNR
542.01 1 14.74 1.632 11.56 14067 1359
553.02 2 10.02 11.923 2.29 681 695 TOI 553.02; active TFOP PC; potential multi
553.03 2 10.02 40.891 2.71 945 461 period updated to 40.89 days with spoc multisector; period could be half; potential multi
555.01 1 14.92 1.942 11.76 32945 857 HATS-76 b
569.01 1 9.43 6.556 8.4 2970 879 crowded field; confirmed BD
570.01 1 9.98 1.469 5.2 1475 1455 V-shaped; crowded field
579.01 1 9.67 1.684 14.07 912 1220 multi year spoc shows no centroid offset but hint of odd-even; host is evolved
580.01 1 9.55 1.55 13.39 3594 5565 variable host; possible secondary not at half phase; multiple stars in pixel; v-shaped
583.01 1 9.38 6.56 11.2 4049 1081 v-shaped
588.01 1 7.31 39.472 16.35 7758 1015 likely too large; possible secondary not at half phase
601.01 1 9.92 3.478 11.33 1177 1229 evolved host
602.01 1 10.05 11.1 28.12 4990 2010 TFOP APC; retired as SB1
609.01 1 9.84 4.97 21.41 12601 908 TFOP APC/SB2
612.01 1 10.19 0.914 5.58 1368 2409 centroid offset not on a particular star; v-shaped
616.01 1 9.98 2.806 0.0 3350 0 transiting object might be too big; no stellar radius; Gaia Rs = 4.24 Rsun would make Rpsimilar-$\sim$20 Re
618.01 1 10.13 7.761 7.86 3110 1250
622.01 1 8.54 6.403 9.3 3815 1274 TOI-622 b
626.01 1 9.51 4.401 19.74 6903 2177 KELT-25 b (Rodriguez Martinez et al; arxiv:1912.01017.pdf)
637.01 1 8.82 2.858 0.0 529 1284 v-shaped; no stellar parameters but possible G/K dwarf host; possible offset to E
640.01 1 10.04 5.004 19.78 7390 1622 TOI-640 b
641.01 1 9.9 1.893 3.12 630 1333
666.01 1 8.79 7.476 14.54 1519 1629 v-shaped; TFOP APC/VPC?/SM
678.01 1 10.87 11.33 3.41 821 808 crowded field
681.01 1 10.66 15.778 18.29 6806 1097 Likely EB; likely too large to be a planet at this irradiation
686.01 1 9.92 3.514 3.34 334 1635
695.01 1 10.73 21.376 2.68 786 563 Was previously designated as EB/AFA; reinstated as PC
696.01 3 10.51 0.86 0.97 720 830 LHS 1678 b
696.02 3 10.51 3.694 1.04 833 511 LHS 1678 c
696.03 3 10.51 4.965 0.86 633 463 LHS 1678 d
697.01 1 9.3 8.608 2.62 555 840 potential L1 planet
699.01 3 9.8 14.801 2.45 345 761 potential L1 planet; possible multi; TOI 699.01
699.02 3 9.8 33.634 2.31 384 636 potential L1 planet; possible multi; TOI 699.02
699.03 3 9.8 672.637 2.97 423 257 could be variability; low SNR; potential multi; only two transits with maximum period of $\sim$672 days; actual period likely shorter
700.01 4 10.88 16.051 2.59 2769 328 TOI-700 c / TOI-700.01
700.02 4 10.88 37.424 1.33 738 247 TOI-700 d / TOI-700.02
700.03 4 10.88 9.977 0.96 460 384 TOI-700 b / TOI-700.03
700.04 4 10.88 27.81 0.95 407 273 TOI-700 e / TOI-700.04
702.01 1 11.75 3.568 2.05 907 676 L1 candidate
703.01 2 10.18 8.673 2.28 725 834 potential L1 candidate in multi-planet system; Gaia Rs=0.91+/-0.04
703.02 2 10.18 22.568 2.48 870 606 possible multi
704.01 1 10.09 3.814 1.11 458 578 LHS 1815 b
706.01 1 8.34 719.038 16.63 4851 245 Likely too large to be a planet at allowed insolations. Two transits two years apart. Actual period is likely shorter.
707.01 2 10.06 52.799 2.4 541 464 potential multi; close to 3:1 resonance
707.02 2 10.06 17.476 1.49 163 671 close to 3:1 resonance; low SNR
709.01 2 14.22 32.377 3.68 52580 2129 TFOP work in progress.
709.02 2 14.22 777.052 1.34 7212 738 1.3 Re candidate around a white dwarf; potential multi-planet system; period likely shorter than maximum period of $\sim$777 days
711.01 2 11.68 18.384 2.15 1010 458 possible multi; potential L1 candidate
711.02 2 11.68 35.943 2.74 1179 366 low SNR; possible multi
712.01 3 9.84 9.531 2.25 916 584 TOI-712 b
712.02 3 9.84 51.699 2.46 1339 357 TOI-712 c
712.04 3 9.84 84.839 2.04 1020 292 TOI-712 d
713.01 2 10.64 35.999 2.23 1087 356 potential L1 planet; possible multi
713.02 2 10.64 1.872 1.06 256 953 potential multi
714.01 2 11.5 4.324 1.5 977 563 potential multi planet system
714.02 2 11.5 10.178 1.57 943 423 possible multi
721.01 1 10.16 12.288 2.09 315 935
722.01 1 11.94 15.3 2.4 1265 519 potential L1 planet; stellar radius is from SPOC Y1 multisector; TIC 38509907 has been joined with TIC 684936227; possible additional candidate at $\sim$83.78 days
723.01 1 10.01 1.418 1.11 204 1284 potential L1 planet; variable host
724.01 1 9.73 3.213 2.06 510 1085 slightly v-shaped; variable host
738.01 1 12.68 3.441 8.9 8647 1089
739.01 1 11.54 9.015 12.23 4800 945 V-shaped
746.01 1 11.38 10.98 10.18 8276 781
785.01 1 11.47 18.626 1.18 425 367 bad transit shape; consistent centroid offset
786.01 2 9.8 12.669 2.31 321 1356 possible multi
786.02 2 9.8 38.554 2.14 329 596 potential multi; centroids scattered between sectors but majority on target
787.01 1 9.47 2.126 1.54 150 1774 centroid variation between sectors; slight depth aperture correlation
788.01 1 8.96 6.487 3.26 113 1048 low S/N; crowded field; possible offset toward faint neighbor TIC 765721735; some odd-even clearer in SPOC TEC plots
790.01 3 8.98 199.578 6.99 1767 423 potential multi
790.02 3 8.98 41.018 2.4 229 717 possible multi near 5:1 resonance
790.03 3 8.98 0.0 6.44 1628 411 single transit exterior to 2 TOIs
794.01 1 9.98 52.407 3.31 372 645
795.01 1 10.22 8.761 2.93 480 1051 centroid variation among sectors; L1 candidate
798.01 1 11.91 95.239 0.0 5934 399 Gaia Teff + astrometry seem to indicate host is an M dwarf; no stellar radius
799.01 1 10.54 5.545 1.94 439 959 L1 candidate
800.01 1 12.25 0.966 3.27 997 1844 blue supergiant host
802.01 1 7.48 3.694 0.65 75 943 Centroid variation among sectors
804.01 1 10.71 2.838 0.0 190 1622 Crowded field. Low SNR. Some depth aperture correlation. Asymmetric transit shape. No stellar radius.
805.01 1 13.16 4.118 1.28 1948 441
807.01 1 11.22 5.27 1.68 512 739 SPOC Y1+Y3 indicates 5-sigma centroid offset; very crowded field
809.01 1 10.27 183.459 3.38 645 416 low S/N
810.01 2 10.47 28.306 2.75 357 696 potential multi
810.02 2 10.47 90.848 2.62 314 472 outter candidate in possible multi planet system
811.01 1 10.93 25.176 14.2 12570 617 could be on companion 100757804; confirmed BD (Carmichael+21)
812.01 1 10.96 13.867 14.39 7930 1131 could be on 363914760
813.01 1 9.82 83.896 7.22 1218 595 TOI-813 b; CTOI from Planet Hunters TESS
841.01 1 12.48 3.493 11.47 14230 914
842.01 1 12.54 2.786 12.0 11629 1196
858.01 1 10.64 3.28 13.81 11048 1402 may also be on 198008002
859.01 1 11.96 5.154 4.6 2522 816
862.01 1 8.99 3.31 0.0 5092 1223 variable host; visual binary separated by $\sim$1 arc sec; no stellar radius; Gaia DR2 Rstarsimilar-$\sim$2.7 and Rp/Rstarsimilar-$\sim$0.07 consistent with possible inflated hot Jupiter
863.01 1 9.92 0.525 1.44 277 2038 low SNR; some scatter in centroids between sectors
864.01 1 11.47 0.521 1.0 627 1006
867.01 1 10.4 15.404 2.29 282 808
870.01 1 10.88 22.038 2.34 1208 362 weak secondary
871.01 1 9.7 14.363 1.69 542 565 some odd-even
872.01 2 11.76 2.24 2.75 1153 1062 L1 candidate
872.02 2 11.76 4.973 2.61 826 814 candidate 2 in potential multi-planet system
873.01 1 12.06 5.931 1.8 1022 518
874.01 1 8.75 5.905 2.01 225 1413 potential L1 planet; possible unrelated transit near TBJD 1512
875.01 1 11.29 11.02 2.34 1077 563
877.01 1 10.58 6.016 13.63 5342 1526 TFOP APC
899.01 1 11.53 12.846 11.89 8527 822
919.01 1 11.18 20.027 23.65 17516 689 large; v-shaped; likely EB; TFOP APC; formerly FP/EB; large radius for orbital period
921.01 1 11.26 5.129 3.38 809 951
924.01 1 11.36 12.127 20.32 19172 776 v-shaped transit and large Rp for given period; Gaia DR2 error 5.51 km/s; likely EB
932.01 2 10.74 19.311 2.75 1223 475 possible multi
932.02 2 10.74 7.986 1.71 460 675 inner candidate in potential multi
933.01 1 10.64 88.935 12.65 4026 526 possible odd-even
934.01 1 12.62 3.782 12.44 18871 952
945.01 1 11.12 0.93 4.6 720 2234
970.01 1 10.52 4.986 3.35 970 1252 some centroid offset not centered on another star
1009.01 1 8.43 1.96 0.0 1708 1456 no stellar radius; Simbad lists as B3III/IV; planet likely larger; odd-even; hot star with extremely rapid rotation
1011.01 1 8.24 2.47 1.45 250 1364
1019.01 1 10.64 5.234 24.29 20797 1575 Large for its insolation
1085.01 1 11.6 3.254 14.95 13620 0 TFOP APC/SB2; 1000pm secondary; phased
1207.01 1 9.22 2.628 0.98 75 1418 centroid offsets scattered around target; sub-Earth Rp
1209.01 1 9.59 40.72 3.24 356 668 Crowded field. Slight depth aperture correlation without offset in difference image. Scattered light in some sectors likely affecting centroid. Asymmetric transit shape.
1212.01 1 10.35 105.585 2.5 463 430 low SNR; close to momentum dumps
1218.01 1 10.34 13.774 1.9 338 730 Crowded field
1219.01 1 10.1 1.914 2.04 248 2462 many fainter stars nearby; v-shaped; slight depth aperture correlation; possible secondary and phase modulations
1221.01 1 10.02 91.683 3.24 713 399 TOI-1221 b
1222.01 1 10.27 10.194 2.74 317 994 TFOP work in progress; epoch is from qlp-s39-tois
1225.01 2 11.7 13.896 2.46 749 686 L1 candidate
1225.02 2 11.7 8.121 2.32 595 821 potential multi; low MES
1226.01 1 11.49 3.929 2.33 486 1123 L1 candidate
1228.01 1 9.08 29.05 2.85 318 1115 HD 56414 b
1232.01 1 11.71 14.255 10.7 7900 786
1338.01 1 11.45 95.174 7.85 2600 255 period confirmed 95.2 days (Veselin B. Kostov et al 2020); TOI-1338 b; circumbinary planet
1406.01 1 11.43 10.574 9.44 5460 1007 Confirmed brown dwarf; paper led by Theron Carmichael
1881.01 1 10.28 1.12 5.96 1901 2267 V-shaped; some odd-even; crowded field; CTOI from Luke Bouma
1937.01 1 12.49 0.947 14.67 16490 1646 TOI-1937 A b
2001.01 1 8.91 2.743 24.42 1518 3128 evolved host
2184.01 1 11.34 6.907 9.36 900 1491 TOI-2184 b
2200.01 1 12.24 2.936 7.19 6790 1147 TFOP work in progress
2205.01 1 15.08 0.889 14.25 41422 1044 near neighbor TIC 737476974; V-shaped; large PC
2208.01 1 15.74 0.674 12.47 103652 804 V-shaped; some odd-even
2218.01 1 11.47 0.346 10.19 2800 3243 V-shaped; small a/Rs; possibly EB
2220.01 1 11.56 1.393 13.36 7098 1793 consistent very small odd-even; possible synchronization in early Y1; V-shaped; large
$2$.01 1 12.68 2.246 17.51 12600 1943
2238.01 1 10.72 3.39 2.41 260 1464 Low-SNR candidate. Poor transit shape but good BLS signal. Slight depth aperture correlation.
2248.01 1 10.46 62.162 10.34 3710 500 Gaia DR2 RV error 6.6 km/s; some phased scatter
2310.01 1 12.71 2.943 13.65 12560 1186 good transit shape; 0.6874 arcsec companion
2329.01 1 11.74 6.535 10.78 6370 925 ephemeris match to nearby neighbor 238920875; this is the brighter star; CTOI from Marco Montalto
2333.01 1 12.35 14.22 8.79 2390 1067 Depth-aperture but no bright stars nearby; CTOI from Marco Montalto
2338.01 1 11.7 22.654 12.28 11530 537 TOI-2338 b; CTOI from Marco Montalto
2339.01 1 11.95 9.773 12.02 2280 1298 CTOI from Marco Montalto
2352.01 1 11.09 1.68 5.37 2020 1496 v-shaped; CTOI from Marco Montalto
2353.01 1 11.9 5.752 6.23 702 1534 low SNR; CTOI from Marco Montalto
2356.01 1 11.71 3.663 3.53 970 1473 Potential L1 candidate; small depth aperture correlation; CTOI from Marco Montalto
2363.01 1 12.1 5.544 13.57 5400 1434 Crowded field; CTOI from Marco Montalto
2368.01 1 11.7 5.175 10.77 13300 810 Slight odd-even and inconsistent transit shapes which can be due to few transits available; CTOI from Marco Montalto
2388.01 1 8.12 1.344 0.0 44 1652 multiple star system; no gaia information; likely SV based on sinusoidal shape in spoc-s1-s65 and multiple signals found at similar period
2392.01 3 10.36 5.912 3.29 416 1431 potential multi
2392.02 3 10.36 3.407 2.96 315 2021 potential multi
2392.03 3 10.36 9.17 2.46 186 1210 low SNR; potential multi
2416.01 1 12.4 8.275 11.55 6510 942
2441.01 2 12.83 12.89 2.97 3521 363
2441.02 2 12.83 5.551 2.03 1282 481 low SNR; interior to TOI 2441.01
2447.01 1 10.01 0.0 10.22 8735 719 Single transit. Orbital period is likely ¿ 30 days.
2448.02 1 10.0 19.758 3.81 483 1035 some depth aperture correlation and possible offset
2449.01 1 9.9 16.654 11.91 10168 1215 single transit
2459.01 1 9.4 19.105 2.94 1906 436 potential L1 candidate; deep transit around $\sim$2200 TBJD in SPOC multisector that may be another planet
2463.01 1 9.03 10.608 11.66 1234 1592 evolved and variable star
2479.01 1 9.31 36.838 3.79 965 566 TFOP work in progress.
2490.01 1 11.28 0.0 7.89 4408 770 CTOI from Rafael Brahm; single at TBJD $\sim$2181
2496.01 1 14.2 1.575 22.96 77504 742 deep transit but Rp ¡ 25 Re; possibly synchronized; likely EB
2501.01 1 11.68 5.326 2.64 1502 642 found in faint-star QLP search; crowded field
2506.01 1 9.93 14.97 3.05 430 0 weak BLS; some variability in raw LC
2507.01 1 9.5 17.305 2.77 657 677 slight odd-even may be from detrending
2518.01 1 10.23 6.898 2.19 550 1095 potential L1 planet; possible centroid offset; slight depth aperture correlation
2519.01 1 10.28 6.955 2.29 770 691
2525.01 2 13.4 23.342 9.73 6474 513 CTOI from Planet Hunters; TOI-2525 b/.01
2525.02 2 13.4 49.245 10.37 15606 400 CTOI from Planet Hunters; TOI-2525 c/.02
2527.01 1 11.17 2.302 12.15 1260 2270 possibly synchronized with variability; CTOI from Samuel Grunblatt
2529.01 1 10.67 0.0 10.97 4240 808 single transit; very slight depth-aperture; CTOI from Rafael Brahm
2531.01 1 12.43 8.148 10.14 13498 790 CTOI from Rafael Brahm; also found in faint-star QLP search
2536.01 1 12.7 31.992 0.0 13054 574 Gaia EDR3 parallax is .92 mas; CTOI from Rafael Brahm; no stellar radius
2539.01 1 11.48 0.986 5.2 1343 2119 SG1 should clear two neighbors
2540.01 2 8.27 12.719 2.59 1144 477 CTOI from George Zhou; inner candidate in potential multi; likely multi
2540.02 2 8.27 22.082 2.29 1019 435 CTOI from George Zhou; outer candidate in potential multi
2541.01 1 9.88 1.412 8.15 520 2858 Weak odd-even difference. Apparent secondary eclipse potentially due to stellar variability.
2544.01 1 10.69 21.069 15.36 3393 953 Eccentric EB with secondaries at 0.54 phase; originally alerted on secondaries at period of 10.5 days; retired as TFOP FP/EB
2549.01 1 11.73 4.858 0.0 11067 1546 no radius; found in faint-star QLP search
2553.01 1 11.78 8.385 11.33 12010 1066
2557.01 1 11.18 9.831 24.45 4990 1595 large Rp
2577.01 1 11.6 3.749 13.25 4860 1635 CTOI from Olmschenk (2021)
2588.01 1 10.64 3.587 10.01 3120 1016 v-shaped
2589.01 1 10.67 61.628 11.74 8878 601 TOI-2589 b
2606.01 1 11.78 1.451 20.54 7270 1944 V-shaped; Potential synchronisation in variability; hint of secondary in multiyear report
2612.01 1 13.23 13.357 19.3 29950 988 found in faint-star QLP search
2617.01 1 13.11 6.319 16.87 28370 958 found in faint-star QLP search
2618.01 1 12.05 8.25 11.87 2250 1764 found in faint-star QLP search; slight odd-even
2622.01 1 13.51 1.221 17.47 25580 1915 found in faint-star QLP search
2624.01 1 13.44 3.024 13.18 26980 1052 found in faint-star QLP search
2625.01 1 13.23 2.148 0.0 22125 1413 found in faint-star QLP search; no stellar radius; phase modulations and secondary in SPOC s1-s69
2626.01 1 12.61 3.456 10.56 20449 847 found in faint-star QLP search
2628.01 1 13.3 0.872 10.55 6610 2291 found in faint-star QLP search
2630.01 1 11.7 9.713 12.32 3540 1027 found in faint-star QLP search; Rp with tic8 almost half of previous Rp ($\sim$22 Re -¿ $\sim$12 Re)
2637.01 1 15.84 0.152 0.0 105147 2710 Host is hot subdwarf; no stellar radius; period likely too short for companion to be planetary; period updated to $\sim$0.15 d with clear additional events at half period in recent SPOC sectors
2645.01 1 12.4 1.444 20.67 43370 1552 found in faint-star QLP search
2646.01 1 13.19 0.313 8.1 4540 2247 found in faint-star QLP search; short period; possible SV
2650.01 1 13.44 2.23 12.78 6708 1652 found in faint-star QLP search
2652.01 1 13.31 2.485 0.0 7950 1583 no radius; found in faint-star QLP search
2653.01 1 12.14 0.927 0.0 1640 1663 no stellar radius; found in faint-star QLP search
2660.01 1 12.08 5.018 10.76 2500 1030 found in faint-star QLP search
2661.01 1 13.05 3.626 12.21 10450 1228 found in faint-star QLP search
2678.01 1 13.44 3.452 10.27 13260 957 found in faint-star QLP search
2683.01 1 12.44 1.529 12.69 2800 1392 slightly V-shaped; found in faint-star QLP search
2684.01 1 13.3 2.476 5.26 4370 1179 found in faint-star QLP search
2685.01 1 13.32 3.471 11.18 10070 972 found in faint-star QLP search
2695.01 1 13.05 0.557 13.24 3250 1398 V-shaped; found in faint-star QLP search
2698.01 1 13.37 3.271 0.0 9670 929 no radius; found in faint-star QLP search
2699.01 1 13.06 5.826 14.02 12800 809 found in faint-star QLP search
2704.01 1 12.71 3.636 16.43 16700 1557 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2721.01 1 12.85 4.163 11.11 12870 1265 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2722.01 1 11.15 0.709 4.67 800 1612 V-shaped; found in faint-star QLP search
2723.01 1 13.32 3.384 16.22 10730 1533 found in faint-star QLP search
2724.01 1 13.19 1.488 18.31 44400 1349 found in faint-star QLP search
2725.01 1 12.06 0.185 4.11 1390 2554 found in faint-star QLP search
2726.01 1 13.4 3.947 18.75 16080 1250 found in faint-star QLP search; correct host star for ephemeris match on 100589632/TOI 2342
2727.01 1 13.05 0.671 12.24 2330 1904 found in faint-star QLP search
2734.01 1 12.68 4.588 5.59 2040 1299 found in faint-star QLP search
2735.01 1 12.95 1.966 13.23 29670 1033 found in faint-star QLP search
2739.01 1 11.9 3.147 12.75 5640 940 V-shaped; found in faint-star QLP search; variable host
2752.01 1 13.16 4.92 12.64 10380 1134 found in faint-star QLP search
2755.01 1 13.29 3.319 9.77 6490 1121 found in faint-star QLP search
2761.01 1 13.22 6.121 17.63 17560 1445 found in faint-star QLP search
2763.01 1 12.0 3.998 0.0 5420 0 no radius; found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2764.01 1 12.69 2.277 10.86 3620 1479 found in faint-star QLP search
2765.01 1 12.13 1.019 7.1 2230 1955 found in faint-star QLP search
2766.01 1 11.66 7.275 17.43 15010 0 found in faint-star QLP search
2785.01 1 13.1 11.505 0.0 15910 665 no radius; found in faint-star QLP search
2787.01 1 13.42 4.322 12.73 2640 1557 found in faint-star QLP search
2788.01 1 13.42 4.065 8.31 7700 1173 found in faint-star QLP search
2791.01 1 12.76 3.748 14.01 5010 1864 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2792.01 1 12.33 7.119 12.54 3920 749 found in faint-star QLP search
2793.01 1 13.29 0.356 6.23 2320 1977 V-shaped; found in faint-star QLP search
2795.01 1 13.48 3.634 11.46 15330 927 found in faint-star QLP search
2798.01 1 12.8 3.994 19.0 14090 1524 found in faint-star QLP search
2799.01 1 13.18 1.437 16.93 23180 1742 found in faint-star QLP search
2800.01 1 12.18 7.754 9.78 6560 937 found in faint-star QLP search
2801.01 1 13.42 5.414 15.52 8400 1020 found in faint-star QLP search
2803.01 1 12.08 1.962 17.45 20910 1882 found in faint-star QLP search
2813.01 1 12.48 4.565 14.65 8572 1789 found in faint-star QLP search; v-shaped
2814.01 1 11.09 5.479 15.18 9196 1403 found in faint-star QLP search; CTOI from Olmschenk et al
2817.01 1 12.12 6.28 9.97 10240 1071 found in faint-star QLP search
2818.01 1 11.32 4.04 14.23 15110 1515 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2820.01 1 13.28 13.345 12.03 24630 620 found in faint-star QLP search
2821.01 1 12.92 0.286 13.69 1990 1571 V-shaped; found in faint-star QLP search
2824.01 1 13.09 2.209 13.79 4710 2901 found in faint-star QLP search
2825.01 1 13.02 5.949 11.62 7490 1284 found in faint-star QLP search
2832.01 1 12.97 2.689 13.09 10930 1655 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2833.01 2 13.07 4.39 0.0 6550 1113 no radius; found in faint-star QLP search
2833.02 2 13.07 1.895 6.72 4247 1473 no stellar radius; two stars within 1 sigma of centroid
2837.01 1 13.5 0.86 0.0 14559 1916 no stellar parameters; V-shaped; found in faint-star QLP search; TFOP BEB/APC
2841.01 1 12.63 2.166 14.46 5930 3298 found in faint-star QLP search
2846.01 1 13.37 2.679 14.0 10530 1636 found in faint-star QLP search
2847.01 1 12.22 2.649 11.66 16770 1194 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2850.01 1 10.66 6.633 7.47 3190 440 V-shaped; found in faint-star QLP search
2851.01 1 13.09 4.826 14.79 7650 1648 found in faint-star QLP search
2852.01 1 13.4 0.401 10.06 2000 2214 found in faint-star QLP search
2853.01 1 13.46 3.292 21.08 26840 1062 slightly V-shaped; found in faint-star QLP search
2856.01 1 12.7 0.599 11.59 1290 1950 found in faint-star QLP search
2857.01 1 12.46 0.622 8.64 2345 3263 found in faint-star QLP search; possible secondary; period updated to 2x alerted period (0.3 days) with spoc-s61-b0A
2858.01 1 13.36 2.418 0.0 4150 1299 no radius; found in faint-star QLP search
2859.01 1 12.94 3.959 20.25 13190 1505 found in faint-star QLP search
2861.01 1 13.03 6.627 11.13 5105 1202 found in faint-star QLP search
2862.01 1 13.04 11.487 15.52 16360 868 found in faint-star QLP search
2864.01 1 12.82 0.478 0.0 2690 2020 no radius; V-shaped; found in faint-star QLP search; multiple stars in pixel; distant host; period likely too short to be planetary companion
2865.01 1 12.68 0.452 9.83 1660 1751 found in faint-star QLP search
2867.01 1 13.49 5.639 14.2 7320 1308 found in faint-star QLP search
2868.01 1 12.97 9.38 12.15 8600 1103 found in faint-star QLP search
2872.01 1 12.89 5.22 14.75 6040 2058 found in faint-star QLP search
2873.01 1 12.66 1.252 0.0 8690 0 no radius; V-shaped; found in faint-star QLP search
2880.01 1 13.29 4.78 13.19 5250 1808 found in faint-star QLP search
2883.01 1 12.37 3.706 0.0 4710 1572 V-shaped; found in faint-star QLP search
2887.01 1 13.3 4.918 17.01 20620 1320 found in faint-star QLP search
2888.01 1 11.03 4.219 0.0 9198 0 no radius; V-shaped; found in faint-star QLP search
2889.01 1 12.97 3.658 0.0 32300 1512 found in faint-star QLP search; no stellar radius
2890.01 1 12.62 3.196 16.46 10760 1314 found in faint-star QLP search; v-shaped; possible odd-even
2891.01 1 11.75 17.308 17.42 8880 873 found in faint-star QLP search
2892.01 1 13.38 14.95 15.17 8070 950 found in faint-star QLP search
2893.01 1 13.3 10.363 7.37 5880 796 found in faint-star QLP search
2896.01 1 12.8 1.348 18.09 2570 1632 found in faint-star QLP search
2897.01 1 13.34 12.004 7.95 6890 826 found in faint-star QLP search
2898.01 1 13.43 3.613 17.57 9702 1786 found in faint-star QLP search
2903.01 1 12.73 3.595 5.1 3070 1015 found in faint-star QLP search; low SNR
2904.01 1 12.55 7.501 7.05 2550 1339 found in faint-star QLP search
2905.01 1 12.41 11.879 0.0 8580 780 found in faint-star QLP search; possible depth aperture correlation; v-shaped; variable host; host is distant and likely evolved so Rp may be too large to be planetary; multiple stars in pixel
2910.01 1 13.34 3.308 0.0 11856 1223 found in faint-star QLP search
2911.01 1 13.38 1.77 10.95 49803 922 found in faint-star QLP search; likely M-dwarf host; Rp from qlp-s36-tois as spoc is assuming Rstar is solar
2912.01 1 11.38 8.761 11.81 4200 1037 found in faint-star QLP search
2914.01 1 13.21 3.052 15.53 5836 2154 found in faint-star QLP search
2915.01 1 12.52 5.13 0.0 7670 992 found in faint-star QLP search; no stellar radius
2921.01 1 12.88 3.004 13.45 11000 1596 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
2923.01 1 12.85 4.323 12.22 9090 1246 found in faint-star QLP search
2924.01 1 12.15 7.35 10.95 5727 1175 found in faint-star QLP search
2932.01 1 12.64 3.257 15.97 5960 1956 found in faint-star QLP search
2944.01 1 13.31 7.793 25.45 16340 1164 found in faint-star QLP search; $\sim$2000-3000 ppm secondary in SPOC s1-s65 at half phase
2945.01 1 12.95 3.201 15.36 5500 2123 found in faint-star QLP search
2948.01 1 13.18 2.756 14.79 14830 1261 slightly V-shaped; found in faint-star QLP search; possible secondary at half phase; no stellar radius in TIC; Gaia DR2 Rstar $\sim$1.18 Rsun
3002.01 1 13.3 7.797 22.98 23150 603 found in faint-star QLP search
3003.01 1 11.92 1.622 7.13 1250 2204 found in faint-star QLP search; noisy but strong BLS
3006.01 1 11.96 8.613 20.68 12840 1009 found in faint-star QLP search; large Rp for orbital period; ”The Gaia DR3 NSS SB1 list has a 51 km/s signal at same period as planet candidate.”; v-shaped
3007.01 1 13.16 4.277 12.03 5000 1508 found in faint-star QLP search
3009.01 1 12.91 3.365 16.72 7700 2373 found in faint-star QLP search
3010.01 1 11.97 9.554 18.04 18260 933 found in faint-star QLP search
3011.01 1 13.25 1.842 0.0 10500 1478 V-shaped; found in faint-star QLP search; no stellar radius; $\sim$1300 ppm secondary in SPOC s1-s65; Gaia DR3 RV error $\sim$7.3 km/s; parallax and temperature and log g suggest host star is distant evolved F star
3013.01 1 12.7 5.472 11.57 1800 1619 found in faint-star QLP search; low SNR; some depth aperture correlation
3014.01 1 13.26 1.795 8.71 2860 2288 found in faint-star QLP search
3015.01 1 13.24 4.275 10.75 3620 1625 found in faint-star QLP search
3016.01 1 13.18 5.114 12.99 7090 1069 found in faint-star QLP search; slight depth aperture correlation
3017.01 1 12.54 3.053 16.92 14090 1727 found in faint-star QLP search
3019.01 1 13.48 5.959 12.8 15090 891 V-shaped; found in faint-star QLP search
3020.01 1 13.44 3.033 19.69 12000 1608 found in faint-star QLP search
3021.01 1 13.56 11.265 0.0 10110 1055 found in faint-star QLP search; no stellar radius
3348.01 1 11.4 2.616 4.05 2340 1614 found in faint-star QLP search
3354.01 1 10.96 46.381 5.75 1108 696
3355.01 1 10.82 15.66 2.14 400 904 low SNR
3360.01 1 10.69 16.738 16.01 2570 521
3363.01 1 10.15 9.743 0.0 1566 0 B9 host; possibly a large planet
3368.01 1 12.83 7.261 15.44 3980 1628 found in faint-star QLP search
3370.01 1 12.02 2.351 6.4 1940 1832 found in faint-star QLP search; multiple stars in pixel
3374.01 1 13.1 2.654 23.74 18470 2022 found in faint-star QLP search
3376.01 1 13.42 4.726 14.79 6910 2218 potential weak secondary at phase 0.5; found in faint-star QLP search
3378.01 1 11.97 2.61 15.94 2844 2540 found in faint-star QLP search
3386.01 1 12.48 0.886 5.05 1900 1811 found in faint-star QLP search
3391.01 1 13.51 4.713 19.76 10493 1643 found in faint-star QLP search; Gaia DR2 Rstar $\sim$2.23; large Rp
3393.01 1 13.62 5.206 11.13 9210 1139 found in faint-star QLP search
3394.01 1 12.14 7.837 7.62 2880 1113 found in faint-star QLP search; possible centroid offset to S
3395.01 1 12.68 3.273 0.0 4030 1069 found in faint-star QLP search
3403.01 1 11.97 11.221 17.09 10721 1032 found in faint-star QLP search
3407.01 1 13.06 1.385 11.98 4150 2006 found in faint-star QLP search
3410.01 1 11.58 3.979 0.0 4000 0 found in faint-star QLP search
3411.01 1 13.14 7.063 0.0 13490 0 found in faint-star QLP search
3413.01 1 12.59 4.711 15.65 5410 1588 V-shaped; found in faint-star QLP search
3414.01 1 11.89 9.736 9.29 2590 1250 found in faint-star QLP search
3429.01 1 12.45 13.931 14.02 6720 806 found in faint-star QLP search
3432.01 1 13.79 9.148 0.0 22476 871 V-shaped; found in faint-star QLP search; no stellar parameters; Gaia DR2 color and parallax suggest host is a red giant
3437.01 1 13.07 4.987 12.45 16931 1039 found in faint-star QLP search
4180.01 1 8.26 1.595 4.32 474 2578 possible odd-even or secondary; variable star
4189.01 1 8.76 46.962 2.47 758 468
4212.01 1 11.97 2.349 0.0 3070 1987 found in faint-star QLP search
4223.01 1 12.8 7.265 13.01 13930 1022 found in faint-star QLP search
4224.01 1 12.28 1.161 10.76 4320 2501 found in faint-star QLP search; v-shaped; possible centroid offset to E
4231.01 1 12.36 3.297 13.46 6550 1885 found in faint-star QLP search
4233.01 1 13.57 1.82 16.63 7330 1898 found in faint-star QLP search
4234.01 1 13.54 0.908 8.54 5020 2617 found in faint-star QLP search
4242.01 1 12.47 3.058 14.25 7765 1613 found in faint-star QLP search; CTOI from Olmschenk et al. (2021)
4243.01 1 10.75 1.089 8.71 412 3762 found in faint-star QLP search
4244.01 1 12.4 1.867 10.21 3606 2050 found in faint-star QLP search
4245.01 1 11.61 3.153 5.72 4103 1076 found in faint-star QLP search; multiple stars in pixel
4246.01 1 12.9 12.886 12.1 11435 848 found in faint-star QLP search
4247.01 1 13.56 4.354 23.03 41586 1176 found in faint-star QLP search
4249.01 1 13.55 6.773 19.13 8606 963 V-shaped; found in faint-star QLP search
4250.01 1 13.42 4.019 13.96 6230 1897 found in faint-star QLP search
4252.01 1 12.66 2.609 13.9 10432 1694 found in faint-star QLP search
4253.01 1 12.9 3.419 0.0 19415 1210 found in faint-star QLP search
4254.01 1 12.54 1.146 14.12 4058 2697 found in faint-star QLP search
4255.01 1 12.52 2.506 12.97 3170 1956 found in faint-star QLP search; possible centroid centered near TIC 82302236
4259.01 1 11.86 8.961 17.09 11334 1000 V-shaped; found in faint-star QLP search
4262.01 1 11.19 5.998 11.36 7349 1234 found in faint-star QLP search
4265.01 1 12.12 3.735 8.62 2360 1085 found in faint-star QLP search
4286.01 1 13.52 1.027 3.28 1430 1925 found in faint-star QLP search
4288.01 1 13.18 6.06 8.07 8270 603 found in faint-star QLP search
4303.01 1 9.46 8.611 4.31 543 1320
4313.01 1 9.6 6.823 3.6 226 1370 L1 candidate
4319.01 1 10.82 0.0 6.55 6743 379 alert as single transit on second event
4322.01 1 10.16 13.419 2.42 348 851 weak signal; possibly off-target; odd transits might not be real
4328.01 1 7.77 703.794 1.86 602 190 two transits; actual period likely shorter than 704 days
4343.01 1 11.26 11.542 1.88 416 791
4344.01 1 9.48 731.481 3.25 245 291 check neighbor TIC 6649664; low MES; possibly a single transit at $\sim$1507 TBJD
4348.01 1 8.4 602.245 2.87 220 305 period may be half the max period of $\sim$602 days
4351.01 1 11.02 20.146 2.44 576 589
4355.01 1 7.23 0.0 3.07 186 291 low SNR; updated to single transit at $\sim$1507.586 TBJD base don s1-s69 not finding original second event at TBJD $\sim$2181
4356.01 1 9.25 579.647 4.05 564 306 period could be half; low SNR; bad transit shape; only two transits
4362.01 1 8.46 7.548 2.83 249 1272 possible odd-even; some bad difference images; possibly SV
4369.01 1 9.33 13.577 3.58 263 1090 two stars in pixel of similar magnitude; L1 candidate; this star is listed as spectroscopic binary with period of 516.83 days (Gaia DR3 NS) and also a visual binary with TIC 278862747
4403.01 1 10.74 4.462 3.04 290 1456 potential L1 planet; 2 sectors show significant centroid offsets not centered on other stars
4404.01 1 10.71 39.624 16.76 4511 673 possible secondary at QLP at phase=0.25 ruled out by later SPOC data
4408.01 1 13.29 2.054 3.51 1994 938
4409.01 1 11.81 92.492 7.77 11488 305
4503.01 1 9.04 710.698 2.09 691 244 two transits; actual period likely shorter than 710.697 days; first event is after momentum dump
4504.01 2 12.55 82.929 11.48 13068 183 8 hour TTV; potential multi
4504.02 2 12.55 2.426 2.66 755 1234 potential multi; L1 candidate
4505.01 1 10.38 19.472 17.33 7063 1052 Gaia DR2 RV error 3.76 km/s; strong stellar pulsations; large PC for period
4507.01 1 10.23 104.616 15.82 6111 562 some odd-even in spoc-s01-s65; may be affected by variable host; possibly too large for insolation
4508.01 1 11.84 3.467 2.11 3657 493 potential L1 planet; low MES
4509.01 1 11.4 0.0 10.49 3178 530 still a single transit as of s69
4510.01 1 10.36 194.243 4.47 811 478 weak signal; just 2 transits
4558.01 1 8.71 6.581 1.68 55 1006 weak signal; evolved host star
4562.01 1 11.53 225.104 13.21 10850 341 TOI-4562 b
4567.01 2 11.92 0.841 1.67 549 1088 potential multi
4567.02 2 11.92 23.542 2.31 1412 358 possible multi; L1 candidate
4663.01 1 13.0 5.292 10.13 17870 745 found in faint-star QLP search
4670.01 1 14.87 15.501 10.83 90070 254 found in faint-star QLP search
4676.01 1 13.22 13.292 8.66 7360 652 found in faint-star QLP search
4680.01 1 13.17 17.431 9.82 8180 714 found in faint-star QLP search
4691.01 1 13.36 8.573 10.41 4880 1076 found in faint-star QLP search
4692.01 1 13.28 4.052 10.72 5310 619 found in faint-star QLP search
4706.01 1 11.38 735.982 12.54 9100 0 period may be several times shorter than max period of $\sim$736 days; found in faint-star QLP search
4710.01 1 12.9 1.48 8.37 3090 1909 found in faint-star QLP search
4746.01 1 11.9 7.54 4.39 2920 759 found in faint-star QLP search
4747.01 1 11.81 5.219 3.83 1500 1035 found in faint-star QLP search
4755.01 1 12.02 22.727 19.08 6510 677 found in faint-star QLP search
4757.01 1 13.5 2.08 11.55 5260 2165 found in faint-star QLP search
4759.01 1 12.26 9.658 9.98 2300 1117 found in faint-star QLP search
4765.01 1 13.05 16.373 0.0 7810 799 found in faint-star QLP search; no stellar radius; Rp be too large to be planetary depending on Rstar
4767.01 1 13.19 2.732 13.24 3880 1927 found in faint-star QLP search
4768.01 1 10.44 7.264 15.08 7988 1295 found in faint-star QLP search
4770.01 1 12.98 1.949 15.83 11670 2165 found in faint-star QLP search; possible offset on TIC 112110893 (Tmagsimilar-$\sim$16)
4774.01 1 12.54 8.6 4.56 1620 1020 found in faint-star QLP search
4775.01 1 11.53 2.473 15.5 2740 1826 found in faint-star QLP search
4778.01 1 12.93 3.807 15.42 8644 1359 found in faint-star QLP search
4779.01 1 13.48 10.749 11.47 9240 639 star marker in diffimage does not look accurate; found in faint-star QLP search
4780.01 1 12.46 4.694 16.49 5310 1546 found in faint-star QLP search
4788.01 1 12.48 4.393 14.03 11110 1134 found in faint-star QLP search
4789.01 1 13.16 7.5 11.25 8750 812 found in faint-star QLP search
4795.01 1 12.92 2.743 0.0 8230 1901 star marker in diffimage does not look accurate; found in faint-star QLP search; no stellar radius
4796.01 1 13.15 3.578 15.32 8730 1415 found in faint-star QLP search
4799.01 1 13.1 4.023 12.89 7560 1496 found in faint-star QLP search
4801.01 1 12.96 3.251 0.0 7880 1348 found in faint-star QLP search; no stellar radius; Rp may be too large to be planetary if Rstar is evolved
4804.01 1 10.23 5.065 8.29 915 1514 found in faint-star QLP search; possible offset to S
4806.01 1 11.75 8.448 4.98 2000 1033 found in faint-star QLP search
4807.01 1 11.97 17.687 8.8 10310 333 V-shaped; found in faint-star QLP search
4812.01 1 12.97 3.314 8.16 3550 1753 found in faint-star QLP search
4816.01 1 13.31 11.247 15.75 10340 920 found in faint-star QLP search
4819.01 1 13.08 1.589 14.03 22530 1245 V-shaped; found in faint-star QLP search
4821.01 1 12.78 5.469 12.67 5940 1137 found in faint-star QLP search
4822.01 1 12.66 4.471 15.89 5140 2415 found in faint-star QLP search
4825.01 1 11.23 2.818 0.0 7916 1290 found in faint-star QLP search; no stellar radius in report; Gaia DR2 has Rstar $\sim$ 2 Rsun which would indicate companion is planetary in size
4828.01 1 13.35 19.029 8.14 8910 536 found in faint-star QLP search
4842.01 1 11.74 5.337 16.55 2780 889 found in faint-star QLP search
4844.01 1 12.54 4.543 11.92 8670 1489 found in faint-star QLP search
4848.01 1 13.1 3.724 18.05 8089 1733 crowded field; found in faint-star QLP search
4851.01 1 9.4 21.609 2.13 320 888 found in faint-star QLP search
4858.01 1 14.81 1.751 16.68 70067 666 found in faint-star QLP search; v-shaped; Rp likely too large for M dwarf host
4879.01 1 13.33 5.376 9.72 4040 1099 found in faint-star QLP search
4888.01 1 12.07 9.419 8.09 2670 1076 found in faint-star QLP search; weak transit shape due to poor detrending
4906.01 1 13.37 11.977 8.82 6760 845 found in faint-star QLP search
4908.01 1 13.22 19.711 10.0 4150 889 found in faint-star QLP search
4923.01 1 12.26 12.918 9.96 4615 805 found in faint-star QLP search
4940.01 1 11.76 25.866 6.76 2613 643 found in faint-star QLP search
4980.01 1 12.93 16.157 0.0 13030 466 found in faint-star QLP search; no stellar radius
4981.01 1 12.48 0.21 11.96 1250 2489 found in faint-star QLP search; short period duration ($\sim$5 hours)
4988.01 1 14.88 2.37 21.17 46129 843 found in faint-star QLP search
5151.01 1 12.83 92.195 16.01 27400 298 source of signal previously alerted as TOI-2239.01 /TIC-388130235 (now retired as FP); this is the real star
5402.01 1 12.38 3.317 1.45 462 810 small star small planet
5541.01 1 12.29 9.018 5.61 3480 892 actual source of signal previous alerted as TOI 2448.01 on TIC 231721005
6093.01 1 10.04 1.271 7.86 1115 2380 V-shaped; possible odd-even
6095.01 1 10.1 2.161 1.87 529 1187 variable host
6098.01 1 8.77 2.73 2.12 362 1048
6100.01 1 11.94 1.228 3.2 1317 1448 Weak signal; short duration; nearby TIC 770748659 (delta Tmag $\sim$ 5) about 1.5 ” away
6102.01 1 13.58 5.371 12.06 14601 1041
6103.01 1 10.67 3.242 13.81 10616 1596 nearby bright star (Tmagsimilar-$\sim$8) 35” away contaminating difference image but signal appears to be on this star
6105.01 1 10.77 3.112 14.38 14170 1204
6107.01 1 11.73 3.79 23.93 7609 2671 large planet around large star
6108.01 1 13.75 3.878 12.53 7843 1330
6264.01 1 10.2 5.693 3.81 291 1529 L1 candidate; possible offset E; slight depth aperture correlation
6271.01 1 13.7 1.604 13.86 19169 1557 multiple stars in pixel; no stellar radius but Gaia DR2 Rs=1.22 so Rpsimilar-$\sim$17 Re
6401.01 1 12.68 3.02 0.0 11319 0 found in faint-star QLP search
6407.01 1 12.12 8.991 10.91 3320 1295 found in faint-star QLP search
6409.01 1 12.24 1.53 19.19 9499 2220 found in faint-star QLP search
6411.01 1 13.08 8.217 7.33 5304 1010 found in faint-star QLP search
6415.01 1 13.42 20.838 10.5 10875 689 found in faint-star QLP search
6416.01 1 13.29 8.233 0.0 29859 0 found in faint-star QLP search
6421.01 1 13.28 3.751 17.42 13942 1527 found in faint-star QLP search
6424.01 1 13.25 2.04 7.5 4429 1863 found in faint-star QLP search
6425.01 1 11.96 12.142 11.8 4584 903 found in faint-star QLP search
6427.01 1 12.46 18.373 8.41 4311 764 found in faint-star QLP search
6428.01 1 13.18 1.676 17.74 13272 1463 found in faint-star QLP search
6435.01 1 11.05 6.116 6.7 3588 754 found in faint-star QLP search
6437.01 1 11.62 8.31 18.26 5667 1603 found in faint-star QLP search
6442.01 1 11.04 2.828 18.52 14846 1611 found in faint-star QLP search
6443.01 1 11.8 9.987 10.29 2884 893 found in faint-star QLP search
6444.01 1 13.34 12.293 14.02 10299 991 found in faint-star QLP search
6448.01 1 12.32 14.844 7.05 7245 641 found in faint-star QLP search
6451.01 1 13.39 5.495 10.08 5504 1296 found in faint-star QLP search
6453.01 1 13.05 6.869 12.82 4404 1506 found in faint-star QLP search
6455.01 1 10.87 0.897 15.85 2703 2622 V-shaped; evolved host
6459.01 1 10.36 5.149 2.77 737 1302 found in faint-star QLP search; possible centroid offset
6471.01 1 13.14 3.076 5.1 2346 1354 found in faint-star QLP search
6472.01 1 13.3 2.681 14.35 6451 2525 found in faint-star QLP search
6475.01 1 12.32 26.599 15.71 8222 1093 found in faint-star QLP search
6480.01 1 11.84 750.24 9.69 6713 135 found in faint-star QLP search; period likely shorter
6483.01 1 12.47 5.208 6.03 4315 570 found in faint-star QLP search
6484.01 1 12.08 18.246 3.87 3157 579 found in faint-star QLP search
6488.01 1 11.82 8.085 3.89 1521 864 found in faint-star QLP search
6493.01 1 12.48 3.067 3.39 3497 687 found in faint-star QLP search
6513.01 1 14.82 2.421 9.94 20899 670 found in faint-star QLP search
6520.01 1 12.72 7.602 19.35 2969 719 found in faint-star QLP search
6521.01 1 14.92 3.581 12.25 41493 661 found in faint-star QLP search
6523.01 1 12.8 7.809 14.05 11239 1065 found in faint-star QLP search
6524.01 1 14.48 0.852 16.07 51018 689 found in faint-star QLP search
6535.01 1 11.74 20.836 7.56 2451 704 found in faint-star QLP search
6549.01 1 12.23 6.108 19.14 14668 1222 Gaia DR3 NSS SB1 list has a 13 km/s signal at P $\sim$7.6 d
6550.01 1 10.81 6.915 15.96 5765 1965 found in faint-star QLP search
6552.01 1 12.39 3.271 13.58 7559 1551
6601.01 1 14.65 32.544 12.0 101822 167 found in faint-star QLP search
6614.01 1 12.57 25.071 11.99 12286 532 found in faint-star QLP search
6618.01 1 11.28 35.789 9.21 2256 796 CTOI from Montalto et al.; found in faint-star QLP search
6621.01 1 12.81 4.281 2.84 2031 778 found in faint-star QLP search
6622.01 1 13.33 37.858 7.36 5180 664 found in faint-star QLP search
6633.01 1 12.18 6.548 2.29 988 878 found in faint-star QLP search
6640.01 1 12.39 16.936 3.6 2618 633 found in faint-star QLP search
6644.01 1 12.57 42.731 4.04 2711 499 found in faint-star QLP search
6654.01 1 10.1 9.516 17.81 487 610 evolved host; Rstar may be smaller ($\sim$5.8 Rsun in Sharma+ 2018); possible depth aperture correlation
6670.01 1 8.89 0.0 9.39 2390 309 single transit; long duration; CTOI from Guoyou Sun
6690.01 2 10.3 761.709 2.56 545 200 two events; real period likely shorter than 761 days; somewhat v-shaped; low MES/SNR
6690.02 2 10.3 904.535 2.73 603 238 alerted as single at TBJD $\sim$1433; now have second transit with max period $\sim$904.534 d; actual period could be shorter
6691.01 1 8.62 0.0 2.95 184 340 single transit; possible additional transits at $\sim$2119 and $\sim$2367 TBJD; still a single transit as of s69
6700.01 1 12.23 15.518 9.48 3717 942 CTOI from Marco Montalto (Montalto et. al 2020)
6701.01 1 11.3 10.018 12.32 1791 1313 v-shaped; some events on/near momentum dumps potentially causing apparent odd-even
6702.01 1 12.98 2.187 3.33 769 960 crowded field; many fainter stars in pixel; v-shaped; possible odd-even
6703.01 1 8.44 49.618 2.47 88 745 low MES/SNR; possible L1 planet; slightly evolved host
6704.01 2 11.53 2.861 0.0 928 1284 potential multi; equally bright neighbor 2 arcsec away; no stellar radius but host is likely K dwarf which would allow for possible L1 Rp
6704.02 2 11.53 6.828 5.09 1559 961 potential multi; likely K dwarf host; equally bright star in same pixel
6707.01 1 12.33 3.835 5.66 1112 1509 possible odd-even
6710.01 1 9.69 2.067 1.51 394 1127 low SNR; variable host
6716.01 1 11.77 4.719 1.01 1735 375 low SNR; flaring host; some transits close to flares
6725.01 1 9.76 2.148 0.0 136 0 no stellar radius; host is likely nearby late K dwarf; Rp likely ¡ 2 Re based on transit depth
6728.01 1 9.61 24.964 4.53 461 768 slightly evolved host star
6813.01 1 12.07 2.356 1.98 740 1150 found in faint-star QLP search
6820.01 1 12.9 2.014 2.55 971 1261 found in faint-star QLP search
6826.01 1 12.18 6.438 2.11 730 879 found in faint-star QLP search
6847.01 1 13.13 1.394 8.87 5131 1545 found in faint-star QLP search; variable host
6851.01 1 12.47 12.01 2.29 1204 513 found in faint-star QLP search
6854.01 1 13.11 2.766 6.36 2476 1258 found in faint-star QLP search
6900.01 1 11.59 13.922 1.72 552 534 low SNR; inconsistent centroids; many fainter stars nearby
6902.01 1 5.68 36.436 2.87 384 705 variable host; interior to HR2562 b
6957.01 1 10.85 160.676 3.22 960 320 some asymmetry in some transits; only 4 events
7047.01 1 11.77 5.273 2.57 1177 1140 L1 candidate; some depth aperture correlation; CTOI from Marco Montalto
Table 5: Continued.
 [^28] The columns give: the TOI identifier of the planet, the multiplicity of the system, the TESS magnitude $T$, the orbital period $P$ in days, the planetary radius $R_{\mathrm{p}}$ in Earth units, the planetary equilibrium temperature in K, and a comment field. All the columns except the second one were copied from the TOI database, 2024-08-20 release (Guerrero et al. 2021)